### v3

# literature.bib

Abstract
We analyze a recently proposed class of algorithms for the problem of sampling from probability distributions $\mu^{\ast}$ in $\mathbb{R}^{d}$ with a Lebesgue density of the form $\mu^{\ast}(x)\propto\exp(-f(Kx)-g(x))$, where $K$ is a linear operator and $f,g$ convex and non-smooth. The method is a generalization of the primal-dual hybrid gradient optimization algorithm to a sampling scheme. We give the iteration’s continuous time limit, a stochastic differential equation in the joint primal-dual variable, and its mean field limit Fokker-Planck equation. Under mild conditions, the scheme converges to a unique stationary state in continuous and discrete time. Contrary to purely primal overdamped Langevin diffusion, the stationary state in continuous time does not have $\mu^{\ast}$ as its primal marginal. Thus, further analysis is carried out to bound the bias induced by the partial dualization, and potentially correct for it in the diffusion. Time discretizations of the diffusion lead to implementable algorithms, but, as is typical in Langevin Monte Carlo methods, introduce further bias. We prove bounds for these discretization errors, which allow to give convergence results relating the produced samples to the target. We demonstrate our findings numerically first on small-scale examples in which we can exactly verify the theoretical results, and subsequently on typical examples of larger scale from Bayesian imaging inverse problems.
Keywords: Langevin sampling, Markov chain Monte Carlo, Fokker-Planck equations, Coupling analysis, Primal-dual algorithms, Bayesian inference, Inverse problems
MSC codes: 35Q84, 47A52, 49N45, 60H10, 62F15, 65J22, 68U10
1 Introduction
We analyze a recently proposed class of algorithms for sampling from a probability distribution on $\mathcal{X}\coloneqq\mathbb{R}^{d}$ with log-concave, non-smooth density. We are interested in target distributions with Lebesgue densities of the form
$\mu^{\ast}(x)=\frac{\exp(-f(Kx)-g(x))}{\int\exp(-f(Kw)-g(w))\,\mathrm{d}w},$ (1.1)
where we assume that the integral in the denominator exists and is finite. The operator $K:\mathcal{X}\to\mathbb{R}^{m}\eqqcolon\mathcal{Y}$ is assumed to be linear and $f:\mathcal{Y}\to\mathbb{R}\cup\{\infty\}$, $g:\mathcal{X}\to\mathbb{R}\cup\{\infty\}$ will be called potential terms and are both assumed convex and lower semi-continuous but not necessarily differentiable. We will denote the full potential by $f\circ K+g\eqqcolon h:\mathcal{X}\to\mathbb{R}\cup\{\infty\}$.
A standard class of algorithms to draw samples in Bayesian inference are Markov chain Monte Carlo (MCMC) methods. Compared to other types of MCMC methods, sampling schemes based on Langevin diffusion processes have proved efficient in inverse problems and imaging applications due to the preferable scaling behaviour to high dimensions [Roberts1996, Durmus2022, Durmus2019]. Typically, MCMC sampling algorithms based on Langevin diffusion perform a time discretisation of the stochastic differential equation (SDE)
$\mathrm{d}X_{t}=\nabla\log\mu^{\ast}(X_{t})\,\mathrm{d}t+\sqrt{2}\,\mathrm{d}W_{t},$ (1.2)
where $W_{t}$ is standard $\mathbb{R}^{d}$-valued Brownian motion. Assuming $\mu^{\ast}$ satisfies a log-Sobolev inequality, it is the unique invariant distribution of the SDE and the law of $X_{t}$ will converge to $\mu^{\ast}$ as $t$ grows. The standard Euler-Maruyama discretization of (1.2) takes the form
$\left\{\begin{array}[]{l}\xi^{n+1}\sim\mathrm{N}(0,I_{d})\\ X^{n+1}=X^{n}+\tau\nabla\log\mu^{\ast}(X_{k})+\sqrt{2\tau}\xi^{n+1},\end{array}\right.$ (1.3)
this iteration is also known as unadjusted Langevin algorithm (ULA). ULA and other time discretizations of (1.2) introduce a bias between the law of the samples and the target $\mu^{\ast}$, hence many works have been concerned with the characterization of the invariant distribution of the Markov chain $(X^{n})_{n}$ and its bias [Roberts1996, Durmus2017, Wibisono2018, Dalalyan2019]. Other typical questions concern preconditioning and convergence speed of the Markov chain and the behaviour of Metropolis-Hastings correction steps to overcome the bias and draw unbiased samples from $\mu^{\ast}$.
In this work, we are concerned with a new type of Langevin sampling algorithm which has been proposed as Unadjusted Langevin Primal-Dual Algorithm (ULPDA) in [Narnhofer2022] and has since been mentioned in [Lau2023, Habring2023]. Unlike ULA or its variants, ULPDA can not be interpreted as a time discretization of (1.2). It rather consists in steps of a primal-dual optimization method, where the same stochastic term $\sqrt{2\tau}\xi$ as in ULA is added to the primal iterate. The iteration takes the form
$\left\{\begin{array}[]{l}\xi^{n+1}\sim\mathrm{N}(0,I_{d})\\ Y^{n+1}=\operatorname{prox}_{\sigma f^{\ast}}\left(Y^{n}+\sigma KX_{\theta}^{n}\right)\\ X^{n+1}=\operatorname{prox}_{\tau g}(X^{n}-\tau K^{T}Y^{n+1})+\sqrt{2\tau}\xi^{n+1}\\ X_{\theta}^{n+1}=X^{n+1}+\theta(X^{n+1}-X^{n}),\end{array}\right.$ (1.4)
where $\tau,\sigma>0$ are step sizes, $\theta>0$ is an overrelaxation parameter, $f^{\ast}$ denotes the convex conjugate of $f$ and the proximal mapping of a convex function $\varphi$ is the unique value $\operatorname{prox}_{\gamma\varphi}(z)\coloneqq(id+\gamma\partial\varphi)^{-1}(z)$.
The motivation for this update comes from the optimization perspective: The ULA iteration (1.3) can be interpreted as gradient descent with an additive stochastic term. While ULA requires to know the gradients of the potentials, the advantage of primal-dual algorithms is that the iteration can be applied without the need to evaluate any gradients or even the proximal mapping of the composition $f\circ K$. $f$ and $g$ need only be convex, not necessarily differentiable. Hence, it appears desirable to obtain more general sampling algorithms by adding stochasticity to other first-order optimization methods like the primal-dual update.
This is, of course, not a rigorous argument for convergence, and until now, to the best of the authors’ knowledge, no convergence theory for ULPDA has been provided. Indeed, the situation seems to be more intricate: When we run ULPDA on a simple toy example, see Figure 1, the distribution of the samples seems to depend more critically on the choice of step size hyperparameters $\tau,\sigma$ of the algorithm than it is known from ULA.
Figure 1: In a simple toy example, we choose $f,g:\mathbb{R}\to\mathbb{R}$ to be quadratic functions, $K>0$ a real number, so that the target $\mu^{\ast}$ is Gaussian with the density depicted in orange. We run ULPDA with very small primal step size $\tau$ and sufficient burn-in period and repeat the experiment for three different ratios $\sigma/\tau$. The gray histograms show the distribution of samples produced by ULPDA. For comparison, the right figure is produced by ULA with the same primal step size $\tau$.
1.1 Related Works
There is a vast literature on unadjusted Langevin algorithms, where we use the term to refer not only to (1.3), but any time discretization of (1.2). The convergence analysis is based on the simple fact that the target distribution is the stationary distribution of the SDE (1.2) and that the dynamics converge to the target in continuous time [Roberts1996]. As an underlying concept, this convergence can be explained by drawing the connection between the diffusion dynamics and, equivalently, the gradient flow of relative entropy with respect to the target in Wasserstein space [Jordan1998, Wibisono2018]. Standard convergence results on ULA typically require a Lipschitz continuous gradient of the potential in order to bound the error introduced by time discretization [Dalalyan2019, Durmus2019, Durmus2019a].
Since non-smooth densities do not satisfy this gradient Lipschitz condition, many works have treated the extension to the non-smooth case. Generally, the approaches can be divided into two classes. The first regularizes non-smooth terms, resulting in a modified, smooth target and applies typical ULA to this modified distribution. The challenge then typically consists in bounding the terms between the target and the modified distribution, and numerically in trading off the introduced bias and convergence speed via the smoothing parameter [Pereyra2016, Durmus2022, Durmus2018, Lau2022, Cai2023]. The second approach consists in finding different discretizations of (1.2), i.e. implicit or semi-implicit schemes, that are implementable and provably convergent. These schemes are usually referred to as proximal or implicit Langevin methods, see e.g. [Pereyra2016, Bernton2018, Salim2020, Klatzer2023, Ehrhardt2023]. A work of particular relevance for our setting is [Habring2023], since the authors assume the same type of splitting of the potential $h=f\circ K+g$ as we do. Their method ‘Prox-Sub’ can be understood as a first-order forward-backward discretization of (1.2), with the forward step in $f\circ K$ realized as a subgradient since $f$ is assumed non-smooth. We will see that this method is intimately connected to the primal-dual sampling algorithm since it forms the limiting case in which the primal-dual stationary distribution can be related to the target. In [Narnhofer2022], the authors proposed the primal-dual scheme that we consider here, in order to generalize from the class of methods that employ purely primal discretizations of (1.2). A partial dualization of the potential is also considered in [Cai2023], but the scheme differs from the one considered here in that it runs an inner iteration to solve the dual optimality condition. Further, the algorithm in [Cai2023] assumes smoothness of $g$ since the updates require evaluation of $\nabla g$. We also refer to [Lau2023] for a recent comparison of current methods applicable in the non-smooth case.
Constraints on the state space can be viewed as a special case of non-smooth potentials, where the constraint is enforced by adding an indicator function to the potential. Several works have considered this special case usually by incorporating a proximal (which amounts to a projection in the case of an indicator) or a reflection step [Brosse2017, Bubeck2018, Melidonis2023]. Mirroring the Langevin dynamics can be another approach to precondition the dynamics as well as to enforce geometrical constraints [Girolami2011, Ahn2021, Li2021, Lau2022].
We also mention that due to the dualization, the primal-dual Langevin sampling strategy can be interpreted as an alternating stochastic descent-ascent scheme. There is a more general interest in the analysis of Langevin descent-ascent schemes in the context of saddle point problems, e.g. for the recovery of Nash equilibria, see [Wang2024]. Our results give answers on the convergence in continuous and discretized time for the case where the saddle point functional is of a Lagrangian form obtained from rewriting a convex optimization problem as a saddle point problem.
1.2 Contributions
Our main contributions are the following (see also Figure 2 for an overview).
  • For the algorithm’s continuous time limit, we prove a contraction result using a coupling argument. This guarantees the dynamics are asymptotically stable and converge to a stationary distribution $\pi_{\lambda}(x,y)$, where $(x,y)\in\mathcal{X}\times\mathcal{Y}\eqqcolon\mathcal{Z}$ in the joint space of primal and dual variable. Crucially, the distribution depends on $\lambda>0$, a constant coupling primal and dual time scales.
  • We show that the marginal of $\pi_{\lambda}$ in the primal variable $x$ is not the target $\mu^{\ast}$. This can be interpreted as being a consequence of $\pi_{\lambda}$ not concentrating its mass on a lower-dimensional submanifold of $\mathcal{X}\times\mathcal{Y}$, although the diffusion coefficient is degenerate with diffusion only acting directly on $x$. In the case of smooth $f^{\ast},g$, we prove this property called hypoellipticity.
  • We consider two potential frameworks for “correcting” the bias between the stationary distribution $\pi_{\lambda}$ and $\pi^{\ast}$ (a concentrated push-forward of the target $\mu^{\ast}$ to $\mathcal{Z}$). Firstly, a proposed modification of the dynamics with non-homogeneous diffusion coefficient can be employed to force $\pi^{\ast}$ to be the stationary solution. The target becomes the stationary distribution, but the discretization of the diffusion SDE is impractical. Secondly, we consider the unmodified scheme, but with the ratio $\lambda$ of dual and primal step sizes in the algorithm growing to infinity. The limit recovers the purely primal overdamped Langevin diffusion. Under regularity assumptions on the dualized potential $f$, we show convergence of $\pi_{\lambda}$ to $\pi^{\ast}$.
  • In discrete time, we show stability of the algorithm in the sense that Wasserstein distances between two different instances are bounded if the potentials are convex. Under stronger regularity assumptions on $f^{\ast},g$, there exists a stationary solution $\pi_{\lambda,\tau}$ to which the distribution of the iterates converges. We then provide a bound of the bias, i.e. a bound on the Wasserstein distance between $\pi_{\lambda}$ and $\pi_{\lambda,\tau}$. This allows distance bounds between the generated samples and $\mu^{\ast}$.
Primal-dual ($0<\lambda<\infty$)Primal ($\lambda=\infty$) Continuous time Contraction to $\pi_{\lambda}$3.4}, Inconsistency {3.5,3.6}, Non-concentration {3.10}, Corrected SDE {3.13}. Stationary solution $\pi^{\ast}=(id,\nabla f\circ K)_{\#}\mu^{\ast}$ with the target $\mu^{\ast}$. Discretized time Stability {4.1}, Contraction to $\pi_{\lambda,\tau}$4.2}. ULA and variants from other discretizations of (1.2). {3.11}{4.8}other works,e.g. [Habring2023] for $h=f\circ K+g$
Figure 2: Overview of the results in this paper. The central theorems are indicated in curled brackets.
The manuscript is organized as follows. In Section 2, we give the considered sampling procedure based on primal-dual optimization and derive its continuous time limit SDE. Since the continuous time limit of the algorithm does not coincide with overdamped Langevin diffusion, we devote Section 3 to the analysis of the continuous time setting and its mean field limit Fokker-Planck equation. The stability and convergence analysis of the algorithm in discrete time is carried out in Section 4. We present numerical illustrations of our most important findings in Section 5 and conclude in Section 6.
2 Primal-Dual Langevin Sampling Schemes
Although the algorithmic approach is not limited to this setting, in the context of Bayesian inference, the target distribution $\mu^{\ast}$ is typically a posterior distribution. Such a posterior target is defined through the likelihood $p(z|x)$ of observing $z$ given the true parameter $x$, combined with a prior distribution $p(x)$. It is given by
$p(x|z)=\frac{1}{Z}p(z|x)p(x),$
where the normalization constant $Z=\int p(z|w)p(w)\,\mathrm{d}w$ is typically intractable in high dimensions. As an example class of problems that we consider, in Bayesian inference for imaging inverse problems the distributions are typically high-dimensional and involve non-smooth terms. This is because model-based priors on image distributions often assume sparse coefficients in some basis, the recovery of which corresponds to minimization of non-smooth functions [Candes2006]. Therefore, many relevant posteriors $p(x|z)$ fall into the class of targets given in (1.1). We start by explaining the motivation for the primal-dual sampling from the point of view of convex, non-smooth optimization.
2.1 Motivation from Primal-Dual Optimization
A common task in Bayesian inverse problems is the computation of the maximum a posteriori estimate (MAP), i.e. the mode of the distribution $\mu^{\ast}=p(\cdot|z)$. This task forms the connection between Bayesian formulations of inverse problems and variational regularization approaches, since it amounts to the problem of computing the solution to the following minimization problem
$\operatorname*{arg\,min}_{x}f(Kx)+g(x).$ (2.1)
In order to solve (2.1), depending on the structure of $f$, $K$ and $g$ it can be helpful to employ techniques from convex duality on parts of the potential $h=f\circ K+g$. A popular approach consists in reformulating the variational problem (2.1) to the saddle point problem
$\min_{x}h(x)$ $=\min_{x}(g(x)+\max_{y}(\langle Kx,y\rangle-f^{*}(y))$
$=\min_{x}\max_{y}s(x,y)$ (2.2)
with $s(x,y):=g(x)+\langle Kx,y\rangle-f^{*}(y)$. Here $f^{\ast}$ denotes the convex conjugate of $f$ defined by
$f^{\ast}(y):=\sup_{x}\left\{\langle x,y\rangle-f(x)\right\}.$
Under mild assumptions on $g$ and $f^{\ast}$, one can prove existence of a saddle point $(x^{\ast},y^{\ast})$ (see section 19 in [Bauschke2011] for details) the primal part $x^{\ast}$ of which is a solution to the original problem (2.1). Instead of minimizing $h$, one can then employ primal-dual type algorithms to solve the saddle point problem (2.2). For the context of imaging inverse problems, we refer to [Chambolle2016] for an overview.
The primal-dual optimization method (primal descent and dual ascent with overrelaxation in the primal variable) for (2.2) takes the form
$\left\{\begin{array}[]{l}Y^{n+1}=\operatorname{prox}_{\sigma f^{\ast}}\left(Y^{n}+\sigma KX_{\theta}^{n}\right)\\ X^{n+1}=\operatorname{prox}_{\tau g}(X^{n}-\tau K^{T}Y^{n+1})\\ X_{\theta}^{n+1}=X^{n+1}+\theta(X^{n+1}-X^{n}),\end{array}\right.$ (2.3)
with step sizes $\tau,\sigma$, overrelaxation parameter $\theta$ and an initial guess $(X^{0},Y^{0})$. By convention, it is set $X_{\theta}^{0}=X^{0}$. For convergence theory of the scheme (2.3) and possible accelerations we refer to [Chambolle2011, Chambolle2016].
In order to draw the parallels between computing the MAP of $\mu^{\ast}$ and drawing samples from $\mu^{\ast}$, note that ULA (1.3) can be viewed as a scheme alternating a gradient descent step in $h=-\log\mu^{\ast}$ and a diffusion step. The diffusion is scaled to ensure the samples have the correct covariance up to a bias which is due to the discretization error and is controlled by the step size $\tau$. Similar to ULA, several first order optimization algorithms have been transformed into a sampling algorithm by inserting a scaled diffusion step at some point in the iteration [Lau2023]. The choice of optimization step is fairly flexible as long as the resulting method can be viewed as an alternating scheme combining a descent step (i.e., a discretized step along the gradient flow of a potential energy) with the diffusion (an exact step along the gradient flow of entropy) [Wibisono2018].
It appears natural to take a similar perspective on primal-dual optimization methods and add a diffusion step at some point in the iteration. Such a scheme, in the form of (1.4), was proposed under the name unadjusted Langevin primal-dual algorithm (ULPDA) in [Narnhofer2022]. In this variant of the algorithm, the diffusion step is added only in the primal variable. We will consider both (1.4) as well as the more general scheme
$\left\{\begin{array}[]{l}\xi^{n+1}\sim\mathrm{N}(0,I_{d+m})\\ Y^{n+1}=\operatorname{prox}_{\sigma f^{\ast}}\left(Y^{n}+\sigma KX_{\theta}^{n+1}\right)+B_{Y}\xi^{n+1}\\ X^{n+1}=\operatorname{prox}_{\tau g}(X^{n}-\tau K^{T}Y^{n})+B_{X}\xi^{n+1}\\ X_{\theta}^{n+1}=X^{n+1}+\theta(X^{n+1}-X^{n}),\end{array}\right.$ (2.4)
with general constant coefficients $B_{X}\in\mathbb{R}^{d\times(d+m)}$ and $B_{Y}\in\mathbb{R}^{m\times(d+m)}$ for the stochastic component. Note the increased dimension of the stochastic term compared to (1.4), which is a special case of (2.4) with $B_{X}=(\sqrt{2}I_{d},0)$ and $B_{Y}=0$.
As a further variant of the algorithm, we remark that in (1.4), we add the stochastic term $\sqrt{2\tau}\xi^{n+1}$ after applying the proximal operator. Similar to other proximal Langevin algorithms, it may instead be added inside the argument of the proximal mapping, effectively making the primal update read
$X^{n+1}=\operatorname{prox}_{\tau g}(X^{n}-\tau K^{T}Y^{n+1}+\sqrt{2\tau}\xi^{n+1}).$
Since the continuous time limit of both versions is the same SDE, our results on asymptotic behaviour in continuous time in Section 3 hold for either variant. In the discrete time analysis in Section 4, we will focus on the version (1.4) with diffusion as the outer step. Similar results can be proven for the other variant and, where necessary, we will mention how our steps need to be modified to obtain them.
2.2 Continuous Time Perspective
In the analysis of ULA’s convergence behaviour, it is helpful to view the update step as an Euler-Maruyama time discretization of the overdamped Langevin diffusion (1.2). Convergence results for ULA are based on the fact that the target measure $\mu^{\ast}$ is the unique invariant distribution of the Markov process $X_{t}$.
In order to analyze ULPDA, we want to define a similar time-continuous limit of (1.4). We assume that the step sizes $\tau,\sigma$ are coupled via the constant $\lim_{\tau,\sigma\to 0}\frac{\sigma}{\tau}=\lambda$ where $0<\lambda<\infty$. Taking both step sizes to zero at this ratio, the formal limiting Markov process $(X_{t},Y_{t})$ satisfies the diffusion equation
$\left\{\begin{array}[]{l}\,\mathrm{d}X_{t}\in-(\partial g(X_{t})+K^{T}Y_{t})\,\mathrm{d}t+\sqrt{2}\,\mathrm{d}W_{t}\\ \,\mathrm{d}Y_{t}\in-\lambda(\partial f^{*}(Y_{t})-KX_{t})\,\mathrm{d}t,\end{array}\right.$ (2.5)
which we can write in a joint variable $Z_{t}=(X_{t},Y_{t})$ as
$\,\mathrm{d}Z_{t}\in-A_{\text{{PD}}}(Z_{t})\,\mathrm{d}t+B_{\text{{PD}}}\,\mathrm{d}W_{t}.$ (2.6)
where we introduced the primal-dual coefficients
$A_{\text{{PD}}}(Z_{t})\coloneqq\begin{pmatrix}\partial g(X_{t})+K^{T}Y_{t}\\ -\lambda(KX_{t}+\partial f^{\ast}(Y_{t}))\end{pmatrix},\qquad B_{\text{{PD}}}\coloneqq\begin{pmatrix}\sqrt{2}I_{d}\\ 0\end{pmatrix}.$ (2.7)
Note in particular that the diffusion coefficient $\Sigma_{\text{{PD}}}\coloneqq\frac{1}{2}B_{\text{{PD}}}B_{\text{{PD}}}^{T}$ is only positive semi-definite (or degenerate) with an $m$-dimensional null space, where $m$ is the dimension of the dual variable $Y$. Standard existence and uniqueness analysis for kinetic Fokker-Planck equations with positive definite diffusion coefficient does not hold anymore.
In the more general scheme (2.4), the diffusion equation takes the form
$\,\mathrm{d}Z_{t}\in-A_{\text{{PD}}}(Z_{t})\,\mathrm{d}t+B_{\text{{gPD}}}\,\mathrm{d}W_{t},\qquad B_{\text{{gPD}}}\coloneqq\begin{pmatrix}B_{X}\\ B_{Y}\end{pmatrix}.$ (2.8)
If $B_{\text{{gPD}}}$ (where ‘gPD’ denotes ‘general primal-dual’) is invertible, then $\Sigma_{\text{{gPD}}}\coloneqq\frac{1}{2}B_{\text{{gPD}}}B_{\text{{gPD}}}^{T}$ becomes positive definite, and the corresponding Fokker-Planck equation has a solution [Pavliotis2014]. We will check whether the stationary solution can be related to the density of the target distribution $\mu^{\ast}$.
3 The Stationary Solution in Continuous Time
[…]
$\mathcal{T}(\mu,\tilde{\mu}):=\inf_{\mathcal{M}\in\Gamma(\mu,\tilde{\mu})}\left(\mathbb{E}_{(x,\tilde{x})\sim\mathcal{M}}\left[\lVert x-\tilde{x}\rVert^{2}\right]\right)^{1/2},$ (3.1)
where $\Gamma(\mu,\tilde{\mu})$ denotes the set of all couplings with marginals $\mu$ and $\tilde{\mu}$, i.e.
[…]
3.1 Fokker-Planck Equation and Existence of a Stationary Solution
The contraction result follows an argument from the literature on overdamped Langevin diffusion, see [Fournier2021, Durmus2019]. Since this simpler case is instructive, we repeat it here and then generalize to the setting of an SDE with any strongly monotone operator as drift coefficient.
3.1.1 Contraction of Overdamped Langevin Diffusion via a Coupling Argument
Suppose for now that the potential $h:\mathbb{R}^{d}\to\mathbb{R}$ is differentiable and $\omega_{h}$-strongly convex. We repeat here an instructive calculation that has been carried out, e.g., in [Fournier2021, Durmus2019] to show stability of the overdamped Langevin diffusion process
$dX_{t}=-\nabla h(X_{t})dt+\sqrt{2}dW_{t}.$ (3.2)
The distribution $\mu_{t}$ of the process $X_{t}$ in (3.2), denoted $\mu_{t}=\operatorname{Law}(X_{t})$, satisfies the Fokker-Planck equation (or forward Kolmogorov equation) [Pavliotis2014] given by
$\frac{\partial\mu_{t}}{\partial t}=\nabla_{x}\cdot(\mu_{t}\nabla h(x))+\Delta_{x}\mu_{t}.$
The target distribution $\mu^{\ast}$ is the unique stationary solution of the Fokker-Planck equation. Furthermore, the PDE is contractive in Wasserstein distance, which can be shown using the following coupling argument. We will call two processes coupled if both solve (3.2) with the same Brownian motion. Let $X_{t},\tilde{X}_{t}$ be coupled with $X_{0}\sim\mu_{0}$ and $\tilde{X}_{0}\sim\tilde{\mu}_{0}$. Then the distribution $\mathcal{M}_{t}$ of the joint variable $(X_{t},\tilde{X}_{t})$ satisfies the Fokker-Planck equation
$\frac{\partial\mathcal{M}_{t}}{\partial t}=\nabla_{x}\cdot(\mathcal{M}_{t}\nabla h(x))+\nabla_{\tilde{x}}\cdot(\mathcal{M}_{t}\nabla h(\tilde{x}))+(\nabla_{x}+\nabla_{\tilde{x}})\cdot(\nabla_{x}+\nabla_{\tilde{x}})\mathcal{M}_{t}.$ (3.3)
Note that we can rewrite this using $(\nabla_{x}+\nabla_{\tilde{x}})\cdot(\nabla_{x}+\nabla_{\tilde{x}})=\Delta_{x+\tilde{x}}$ by a coordinate change $(x,\tilde{x})\mapsto(x+\tilde{x},x-\tilde{x})$. For any distance function of the form $d(x,\tilde{x})=\rho(\lVert x-\tilde{x}\rVert)$, we directly see $\Delta_{x+\tilde{x}}d(x,\tilde{x})=0$. Using this and the property (3.3) of the joint distribution $\mathcal{M}_{t}$, a contraction in Wasserstein distance can hence be shown in the following way. We start by estimating
$\frac{1}{2}\frac{d}{dt}\mathbb{E}_{(x,\tilde{x})\sim\mathcal{M}_{t}}\left[\lVert x-\tilde{x}\rVert^{2}\right]$ $=\frac{1}{2}\frac{d}{dt}\int\lVert x-\tilde{x}\rVert^{2}\mathcal{M}_{t}(\mathrm{d}x,\mathrm{d}\tilde{x})$
$=\frac{1}{2}\int\lVert x-\tilde{x}\rVert^{2}\left[\nabla_{x}\cdot(\mathcal{M}_{t}\nabla h(x))+\nabla_{\tilde{x}}\cdot(\mathcal{M}_{t}\nabla h(\tilde{x}))+\Delta_{x+\tilde{x}}\mathcal{M}_{t}\right]\,\mathrm{d}x\,\mathrm{d}\tilde{x}$
$=-\int(x-\tilde{x})\cdot(\nabla h(x)-\nabla h(\tilde{x}))\mathcal{M}_{t}(\mathrm{d}x,\mathrm{d}\tilde{x})$
$\leq-\omega_{h}\mathbb{E}_{(x,\tilde{x})\sim\mathcal{M}_{t}}\left[\lVert x-\tilde{x}\rVert^{2}\right],$
where we used that due to strong convexity of $h$, we have $(x-\tilde{x})\cdot(\nabla h(x)-\nabla h(\tilde{x}))\geq\omega_{h}\lVert x-\tilde{x}\rVert^{2}$. Applying now Gronwall’s lemma, we obtain
$\mathbb{E}_{(x,\tilde{x})\sim\mathcal{M}_{t}}\left[\lVert x-\tilde{x}\rVert^{2}\right]\leq e^{-2\omega_{h}t}\mathbb{E}_{(x,\tilde{x})\sim\mathcal{M}_{0}}\left[\lVert x-\tilde{x}\rVert^{2}\right].$
The left side is bounded from below by $\mathcal{W}_{2}^{2}(\mu_{t},\tilde{\mu}_{t})$ by definition. By taking the infimum over all couplings $\mathcal{M}_{0}$ (with marginals $\mu_{0}$ and $\tilde{\mu}_{0}$) on the right hand side, we obtain
$\mathcal{W}_{2}(\mu_{t},\tilde{\mu}_{t})\leq e^{-\omega_{h}t}\mathcal{W}_{2}(\mu_{0},\tilde{\mu}_{0}).$
In particular, if we choose one of the processes to be initialized at the target distribution as $\tilde{\mu}_{0}=\mu_{\ast}$, then the distribution of $\tilde{X}_{t}$ is stationary at $\tilde{\mu}_{t}=\mu^{\ast}$ and by the above we obtain the convergence result
$\mathcal{W}_{2}(\mu_{t},\mu^{\ast})\leq e^{-\omega_{h}t}\mathcal{W}_{2}(\mu_{0},\mu^{\ast}).$
3.1.2 Contraction and Stationary Solution of the Primal-Dual Diffusion Equation
We can generalize the above argument allowing an analysis of the time-continuous limit (2.6) of the primal-dual sampling scheme (1.4). Although the primal-dual dynamics are not the gradient flow of a corresponding relative entropy anymore, contraction can equivalently be shown for a coupling of two processes. For that, let $Z_{t}$ satisfy a general stochastic differential inclusion with non-homogeneous coefficients
$\mathrm{d}Z_{t}\in-A(Z_{t})\,\mathrm{d}t+B(Z_{t})\,\mathrm{d}W_{t}.$ (3.4)
Here $W_{t}$ is Brownian motion in $\mathbb{R}^{K}$ and $B:\mathbb{R}^{d}\to\mathbb{R}^{d\times K}$. We assume the drift $A$ to be a set-valued monotone operator, meaning
$\langle w-\tilde{w},z-\tilde{z}\rangle\geq 0,\qquad\forall w\in A(z),\tilde{w}\in A(\tilde{z}).$
We say that $A$ is $\omega$-strongly monotone if the following stronger condition is satisfied:
$\langle w-\tilde{w},z-\tilde{z}\rangle\geq\omega\lVert z-\tilde{z}\rVert^{2},\qquad\forall w\in A(z),\tilde{w}\in A(\tilde{z}).$
Proposition 3.1.
Let $Z_{t}$ be a solution to (3.4) and denote by $\pi_{t}$ the density w.r.t. the Lebesgue measure of the distribution of $Z_{t}$. Then $\pi_{t}$ satisfies the Fokker-Planck equation
$\frac{\partial\pi_{t}}{\partial t}=\nabla_{z}\cdot\left(w\pi_{t}+\Sigma(z)\nabla_{z}\pi_{t}\right),\qquad w\in A(z).$ (3.5)
with $\Sigma(z)=\frac{1}{2}B(z)B(z)^{T}$.
Since we are repeatedly going to make use of coupling arguments, note that by applying Proposition 3.1 to a coupled diffusion
$\begin{pmatrix}\mathrm{d}Z_{t}\\ \mathrm{d}\tilde{Z}_{t}\end{pmatrix}\in-\begin{pmatrix}A(Z_{t})\\ A(\tilde{Z}_{t})\end{pmatrix}\mathrm{d}t+\begin{pmatrix}B(Z_{t})\\ B(\tilde{Z}_{t})\end{pmatrix}\mathrm{d}W_{t},$ (3.6)
we can also obtain a Fokker-Planck equation for the joint density of the coupled processes.
Proposition 3.2.
Let $Z_{t},\tilde{Z}_{t}$ be two coupled processes that both solve (3.4) with the same Brownian motion, i.e. $(Z_{t},\tilde{Z}_{t})$ solves (3.6). Then the joint distribution $\Pi_{t}(z,\tilde{z})$ of $(Z_{t},\tilde{Z}_{t})$ satisfies the Fokker-Planck equation
$\frac{\partial\Pi_{t}}{\partial t}=\nabla_{z}\cdot\left(w\pi_{t}\right)+\nabla_{\tilde{z}}\cdot\left(\tilde{w}\Pi_{t}\right)+(\nabla_{z}+\nabla_{\tilde{z}})\cdot(\Sigma(z,\tilde{z})(\nabla_{z}+\nabla_{\tilde{z}})\Pi_{t}),\qquad w\in A(z),\tilde{w}\in A(\tilde{z}).$ (3.7)
For showing the contractivity of overdamped Langevin dynamics in Section 3.1.1, the relevant requirement was the strong convexity of $h$, implying that $\nabla h$ is a strongly monotone operator. The computation can be directly generalized under a strong monotonicity assumption on $A$.
Lemma 3.3.
Suppose $A$ is $\omega$-strongly monotone. Then (3.4) is contractive in the sense that, if $Z_{t},\tilde{Z}_{t}$ are two coupled processes that solve (3.4) with the same Brownian motion, then for $t\geq 0$ their associated distributions $\pi_{t},\tilde{\pi}_{t}$ satisfy
$\mathcal{W}_{2}(\pi_{t},\tilde{\pi}_{t})\leq e^{-\omega t}\mathcal{W}_{2}(\pi_{0},\tilde{\pi}_{0}).$
Proof.
With a change of variable we can write $\nabla_{z}+\nabla_{\tilde{z}}=\nabla_{z+\tilde{z}}$. By employing (3.7), we obtain the same type of inequality as in the case of overdamped Langevin diffusion that
$\frac{1}{2}\frac{\,\mathrm{d}}{\,\mathrm{d}t}\mathbb{E}_{(z,\tilde{z})\sim\Pi_{t}}\left[\lVert z-\tilde{z}\rVert^{2}\right]$ $=\frac{1}{2}\frac{\,\mathrm{d}}{\,\mathrm{d}t}\int\lVert z-\tilde{z}\rVert^{2}\Pi_{t}(\mathrm{d}z,\mathrm{d}\tilde{z})$
$=\frac{1}{2}\int\lVert z-\tilde{z}\rVert^{2}\left(\nabla_{z}\cdot\left(w\Pi_{t}\right)+\nabla_{\tilde{z}}\cdot\left(\tilde{w}\Pi_{t}\right)+(\nabla_{z+\tilde{z}})\cdot(\Sigma(\nabla_{z+\tilde{z}})\Pi_{t})\right)\,\mathrm{d}z\,\mathrm{d}\tilde{z}$
$=-\int\left((z-\tilde{z})\cdot w+(\tilde{z}-z)\cdot\tilde{w}\right)\Pi_{t}(\mathrm{d}z,\mathrm{d}\tilde{z})$
$\leq-\omega\mathbb{E}_{(z,\tilde{z})\sim\Pi_{t}}\left[\lVert z-\tilde{z}\rVert^{2}\right]$
In the second line we used Proposition 3.2 and in the third line partial integration together with the fact that $(\nabla_{z}+\nabla_{\tilde{z}})(\lVert z-\tilde{z}\rVert^{2})=0$. We now apply Gronwall’s Lemma and obtain
$\mathbb{E}_{(z,\tilde{z})\sim\Pi_{t}}\left[\lVert z-\tilde{z}\rVert^{2}\right]\leq e^{-2\omega t}\mathbb{E}_{(z,\tilde{z})\sim\Pi_{0}}\left[\lVert z-\tilde{z}\rVert^{2}\right]$
which implies
$\mathcal{W}_{2}(\pi_{t},\tilde{\pi}_{t})\leq e^{-\omega t}\mathbb{E}_{(z,\tilde{z})\sim\Pi_{0}}\left[\lVert z-\tilde{z}\rVert^{2}\right].$
Taking the infimum over all couplings of $\pi_{0}$ and $\tilde{\pi}_{0}$ on the right hand side concludes the proof. ∎
Consider now the general primal-dual type diffusion (2.8) and denote $\pi_{t}$ for the law of $Z_{t}$. By Proposition 3.1, $\pi_{t}$ solves the Fokker-Planck equation
$\frac{\partial\pi_{t}}{\partial t}=\nabla_{z}\cdot\left(w\pi_{t}+\Sigma_{\text{{gPD}}}\nabla_{z}\pi_{t}\right),\qquad w\in A_{\text{{PD}}}(z).$ (3.8)
with $\Sigma_{\text{{gPD}}}=\frac{1}{2}B_{\text{{gPD}}}B_{\text{{gPD}}}^{T}$ as in (2.8) and $A_{\text{{PD}}}$ given in (2.7). The drift coefficient $A_{\text{{PD}}}$ is a monotone operator.
[…]
The transport distance (3.1) on $\mathcal{P}_{2}(\mathcal{Z})$ induced by $\lVert\cdot\rVert_{a,b}$ will be denoted by $\mathcal{T}_{a,b}$. Note that the equivalence of norms on $\mathbb{R}^{n}$ implies equivalence of all $\mathcal{T}_{a,b}$ for $a,b>0$.
Theorem 3.4.
Assume that $g,f^{\ast}$ are $\omega_{g}$-strongly convex and $\omega_{f^{\ast}}$-strongly convex, respectively, with $\omega_{g},\omega_{f^{\ast}}>0$. Let $Z_{t},\tilde{Z}_{t}$ be two coupled processes that solve the general primal-dual diffusion equation (2.8) with corresponding densities $\pi_{t}$ and $\tilde{\pi}_{t}$. Then it holds
$\mathcal{T}_{1,\lambda^{-1}}(\pi_{t},\tilde{\pi}_{t})\leq e^{-\omega t}\mathcal{T}_{1,\lambda^{-1}}(\pi_{0},\tilde{\pi}_{0}),$
where $\omega=\min\{\omega_{g},\lambda\omega_{f^{\ast}}\}>0$. In particular, the dynamics are contractive and there exists a unique stationary distribution of the continuous time primal-dual diffusion. Making the dependence on $\lambda$ explicit, we denote this stationary distribution by $\pi_{\lambda}$.
Proof.
We employ Lemma 3.3, hence the only part we need to show is the strong monotonicity of the negative drift coefficient $A_{\text{{PD}}}$ defined by
$A_{\text{{PD}}}=\begin{pmatrix}\partial g&K^{T}\\ -\lambda K&\lambda\partial f^{\ast}\end{pmatrix}.$
We write $z=(x,y)\in\mathcal{X}\times\mathcal{X}=\mathcal{Z}$ for the joint primal and dual variable and equip $\mathcal{Z}$ with the weighted inner product $\langle\cdot,\cdot\rangle_{1,\lambda^{-1}}$. Let now $w\in A_{\text{{PD}}}(z)$ and $\tilde{w}\in A_{\text{{PD}}}(\tilde{z})$. Then there are $u\in\partial g(x)$ and $v\in\partial f^{\ast}(y)$ such that $w=(u+K^{T}y,\lambda(v-Kx))$. Equivalently, if $\tilde{w}\in A(\tilde{z})$ then there exist $\tilde{u}\in\partial g(\tilde{x})$ and $\tilde{v}\in\partial f^{\ast}(\tilde{y})$ with $\tilde{w}=(\tilde{u}+K^{T}\tilde{y},\lambda(\tilde{v}-K\tilde{x}))$. It follows that
$\langle w-\tilde{w},z-\tilde{z}\rangle_{1,\lambda^{-1}}$ $=\langle u-\tilde{u},x-\tilde{x}\rangle+\lambda\langle v-\tilde{v},y-\tilde{y}\rangle_{\lambda^{-1}}$
$\qquad+\langle K^{T}(y-\tilde{y}),x-\tilde{x}\rangle-\lambda\langle K(x-\tilde{x}),y-\tilde{y}\rangle_{\lambda^{-1}}$
$=\langle u-\tilde{u},x-\tilde{x}\rangle+\langle v-\tilde{v},y-\tilde{y}\rangle$
$\geq\omega_{g}\lVert x-\tilde{x}\rVert^{2}+\omega_{f^{\ast}}\lVert y-\tilde{y}\rVert^{2}$
$\geq\omega\lVert z-\tilde{z}\rVert_{1,\lambda^{-1}}^{2}.$
The stationary solution of (3.8) will be denoted by $\pi_{\lambda}$. We will also use the $x$-marginal $\mu_{\lambda}$ of $\pi_{\lambda}$, defined as $\mu_{\lambda}(x)\coloneqq\int_{\mathcal{Y}}\pi_{\lambda}(x,y)\,\mathrm{d}y.$
3.2 Non-Concentration of the Stationary Solution
Having established the existence of $\pi_{\lambda}$, we can try to relate the algorithmic, time discretized scheme to it. Start by considering the special case (2.6) where diffusion is only added in the primal variable. While overdamped Langevin dynamics is designed to have the target $\mu^{\ast}$ as its stationary solution, this is not the case in the primal-dual setting. In general, $\mu^{\ast}\neq\mu_{\lambda}$, in other words $\mu^{\ast}$ is not the $x$-marginal of $\pi_{\lambda}$. We demonstrate this on a simple example of a Gaussian target, where we can solve the Fokker-Planck equation (3.8) in closed form.
Theorem 3.5.
Consider a one-dimensional setting with the target $\mu^{\ast}(x)\propto\exp(-h(x))$ given by
$f(y)\coloneqq\frac{1}{2c_{f}}y^{2},\ g(x)\coloneqq\frac{1}{2c_{g}}x^{2},\ h(x)\coloneqq f(kx)+g(x)=\frac{k^{2}c_{g}+c_{f}}{2c_{f}c_{g}}x^{2},\qquad x,y\in\mathbb{R},$
where $k\neq 0$ and $c_{g},c_{f}>0$. It holds $f^{\ast}(y)=\frac{c_{f}}{2}y^{2}$ for all $y\in\mathbb{R}$. The primal-dual diffusion equation (2.6) for the joint variable $Z_{t}=(X_{t},Y_{t})^{T}\in\mathbb{R}^{2}$ reads
$\mathrm{d}Z_{t}=-A_{\text{{PD}}}Z_{t}\mathrm{d}t+B_{\text{{PD}}}\mathrm{d}W_{t},\quad\text{with }A_{\text{{PD}}}=\begin{pmatrix}c_{g}^{-1}&k\\ -\lambda k&\lambda c_{f}\end{pmatrix},B_{\text{{PD}}}=\begin{pmatrix}\sqrt{2}\\ 0\end{pmatrix}$
where $W_{t}$ is one-dimensional Brownian motion and the weight $\lambda>0$ relates primal and dual time scales. Denoting $E(t):=(e_{ij}(t))_{i,j}:=\exp(-tA_{\text{{PD}}}^{T})$, the following statements hold:
  1. (i)
    For any (possibly singular) initialization $\pi_{0}$ such that $Z_{0}\sim\pi_{0}$, the distribution $\pi_{t}\coloneqq\operatorname{Law}(Z_{t})$ has a $C^{\infty}$-smooth Lebesgue density for $t>0$.
  2. (ii)
    If $\pi_{0}$ is Gaussian (including degenerate Gaussians with singular covariance matrix), then $\pi_{t}$ is (non-degenerate) Gaussian for all $t>0$. As $t\to\infty$, the solution $\pi_{t}$ converges to the Gaussian stationary $\pi_{\lambda}$ with mean zero and covariance matrix
    $\frac{1}{(c_{f}+k^{2}c_{g})(1+\lambda c_{f}c_{g})}\begin{pmatrix}c_{g}(c_{f}+k^{2}c_{g})+\lambda c_{f}^{2}c_{g}^{2}&k\lambda c_{f}c_{g}^{2}\\ k\lambda c_{f}c_{g}^{2}&k^{2}\lambda c_{g}^{2}\end{pmatrix}.$
    The $x$-marginal $\mu_{\lambda}$ is not equal to the target $\mu^{\ast}$. However, the limiting joint distribution as $\lambda\to\infty$ is the degenerate Gaussian $(id,f^{\prime}(k\,\cdot))_{\#}\mu^{\ast}$ whose $x$-marginal is the target $\mu^{\ast}$.
Proof.
From the diffusion equation, it follows that the distribution $\pi_{t}(z)=\pi_{t}(x,y)$ of $Z_{t}$ satisfies the Fokker-Planck equation
$\partial_{t}\pi_{t}=\nabla_{z}\cdot(\pi_{t}A_{\text{{PD}}}z)+\Delta_{x}\pi_{t},$
with an initial value $\pi_{0}$. The Cauchy problem can be solved in closed form, note that a similar computation has been carried out in [Hoermander1967]. By taking the Fourier transform in the variable $z$ (denoted $\mathcal{F}$), we obtain the first order initial value problem for $p_{t}(\zeta)=\mathcal{F}(\pi_{t})(\zeta)$ with $\zeta=(\xi,\omega)^{T}$:
$\left\{\begin{array}[]{l}\partial_{t}p_{t}=-\zeta^{T}A_{\text{{PD}}}\nabla_{\zeta}p_{t}-\xi^{2}p_{t}\\ p_{0}(\xi,\omega)=\mathcal{F}\pi_{0}.\end{array}\right.$
The corresponding characteristic equations are
$\dot{\zeta}(t)$ $=A_{\text{{PD}}}^{T}\zeta(t)$
$\dot{p}(t)$ $=-\xi^{2}(t)p(t),$
and they are solved by
$\zeta(t)$ $=\exp(tA_{\text{{PD}}}^{T})\zeta_{0}$
$p(t)$ $=p_{0}(\zeta_{0})\exp\left(-\int_{0}^{t}\xi^{2}(s)\,\mathrm{d}s\right)$
By eliminating the initial value $\zeta_{0}=\exp(-tA_{\text{{PD}}}^{T})\zeta(t)$, the solution for the initial value problem is
$p_{t}(\zeta)=p_{0}(\exp(-tA_{\text{{PD}}}^{T})\zeta)\exp\left(-\frac{1}{2}\int_{0}^{t}(B_{\text{{PD}}}^{T}\exp(-sA_{\text{{PD}}}^{T})\zeta)^{2}\,\mathrm{d}s\right).$
The quadratic form inside the integral is positive semidefinite in the variable $\zeta$. For $t>0$, it is also positive definite: Otherwise, there would need to be a $\zeta$ such that $B_{\text{{PD}}}^{T}\exp(-sA_{\text{{PD}}}^{T})\zeta=0$ for all $s\in[0,t]$, which implies $B_{\text{{PD}}}^{T}(A_{\text{{PD}}}^{T})^{k}\zeta=0$ for all $k\geq 0$. This is equivalent to $(0,1)^{T}$ being an element of an invariant subspace of $A_{\text{{PD}}}^{T}$, which is impossible since the eigenvectors of $A_{\text{{PD}}}^{T}$ are of the form
$\left(\frac{c_{g}^{-1}-\lambda c_{f}}{2k}\pm\frac{1}{2k}\sqrt{(c_{g}^{-1}-\lambda c_{f})^{2}-4\lambda k^{2}},1\right)^{T},$
neither of which is a multiple of $(0,1)^{T}$. From the positive definiteness of the quadratic form it follows that $p_{t}$ is a Schwartz function for every $t>0$. In particular, its inverse Fourier transform $\pi_{t}$ is $C^{\infty}$-smooth in the variable $z$ for $t>0$, which proves (i).
If the initialization $\pi_{0}$ is Gaussian with mean $z_{0}=(x_{0},y_{0})$ and (possibly singular) covariance $\Sigma_{0}$, we can simply take the inverse Fourier transform to obtain the solution $\pi_{t}$. It is worth noting, however, that we do not even have to compute the full distribution $\pi_{t}$ in order to analyze the $x$-marginal $\mu_{t}$ of $\pi_{t}$. Since marginalizing out the dual variable $y$ in $\pi_{t}$ corresponds to evaluating $p_{t}$ at $\zeta=(\xi,0)$ in Fourier space, we get, denoting by $\mathcal{F}^{-1}_{\xi}$ the inverse Fourier transform in $\xi$ and $E(t):=(e_{ij}(t))_{i,j}:=\exp(-tA_{\text{{PD}}}^{T})$,
$\mu_{t}(x)$ $:=\int\pi_{t}(x,y)\,\mathrm{d}y$
$=\mathcal{F}^{-1}_{\xi}(p_{t}(\xi,0))(x)$
$\propto\int\exp\left(i\xi(x-z_{0}\cdot e_{1}(t))-\frac{\xi^{2}}{2}\left(\lVert(e_{1}(t)\rVert_{\Sigma_{0}}^{2}+2\int_{0}^{t}e^{2}_{11}(s)\,\mathrm{d}s\right)\right)\,\mathrm{d}\xi$
$=\exp\left(-\frac{1}{2}(x-z_{0}\cdot e_{1}(t))^{2}\left(\lVert(e_{1}(t)\rVert_{\Sigma_{0}}^{2}+2\int_{0}^{t}e^{2}_{11}(s)\,\mathrm{d}s\right)^{-1}\right),$
where we abbreviated $e_{1}(t)$ for the first column of $E(t)$. Hence, the primal variable follows a Gaussian distribution at every time $t>0$, with mean $z_{0}^{T}e_{1}(t)$ and variance $\lVert(e_{1}(t)\rVert_{\Sigma_{0}}^{2}+2\int_{0}^{t}e^{2}_{11}(s)\,\mathrm{d}s$. Due to the contraction result Theorem 3.4, we know that $\mu_{t}$ converges to the stationary distribution $\mu_{\lambda}$ as $t\to\infty$. Indeed, this holds here since $\lim_{t\to\infty}E(t)=0$, so $\mathbb{E}_{X\sim\mu_{\lambda}}[X]=0$. The variance of $\mu_{\lambda}$ is
$\operatorname{Var}_{X\sim\mu_{\lambda}}[X]=2\int_{0}^{\infty}e^{2}_{11}(s)\,\mathrm{d}s=\frac{c_{g}(c_{f}+k^{2}c_{g})+\lambda c_{f}^{2}c_{g}^{2}}{(c_{f}+k^{2}c_{g})(1+\lambda c_{f}c_{g})},$
which does not coincide with the variance of the target distribution $\mu^{\ast}$. The target is however approximated as $\lambda$ becomes large, since $\lim_{\lambda\to\infty}\operatorname{Var}[\mu_{\lambda}]=c_{g}c_{f}/(c_{f}+k^{2}c_{g})$.
Analogously, the full solution $\pi_{t}$ of the Fokker-Planck equation can be computed by taking the inverse Fourier transform in $\zeta$ above. We obtain that $\pi_{t}$ is Gaussian at every time $t>0$ with mean $E(t)^{T}z_{0}$ and positive definite covariance matrix
$\operatorname{Var}_{Z\sim\pi_{t}}[Z]=E(t)^{T}\Sigma_{0}E(t)+2\begin{pmatrix}\int_{0}^{t}e_{11}^{2}\,\mathrm{d}s&\int_{0}^{t}e_{11}e_{12}\,\mathrm{d}s\\ \int_{0}^{t}e_{11}e_{12}\,\mathrm{d}s&\int_{0}^{t}e_{12}^{2}\,\mathrm{d}s\end{pmatrix}.$
In the limit $t\to\infty$, we obtain the stationary distribution which has mean zero and covariance matrix
$\operatorname{Var}_{Z\sim\pi_{\lambda}}[Z]=\frac{1}{(c_{f}+k^{2}c_{g})(1+\lambda c_{f}c_{g})}\begin{pmatrix}c_{g}(c_{f}+k^{2}c_{g})+\lambda c_{f}^{2}c_{g}^{2}&k\lambda c_{f}c_{g}^{2}\\ k\lambda c_{f}c_{g}^{2}&k^{2}\lambda c_{g}^{2}\end{pmatrix}$
In the limit $\lambda\to\infty$, the stationary distribution approaches a degenerate Gaussian (in the sense that the covariance matrix is singular), with mean zero and covariance matrix
$\frac{1}{(c_{f}+k^{2}c_{g})}\begin{pmatrix}c_{f}c_{g}&kc_{g}\\ kc_{g}&k^{2}c_{f}^{-1}c_{g}\end{pmatrix}.$
This is identical to $(id,f^{\prime}(k\,\cdot))_{\#}\mu^{\ast}$, proving (ii). ∎
The negative result Theorem 3.5 naturally raises the question whether the diffusion equation can be modified such that the target becomes its stationary distribution. While we will consider the more general case of non-homogeneous, problem-dependent diffusion coefficients in Section 3.4, the (algorithmically) most straightforward approach would be to add homogeneous diffusion not only in the primal variable $X^{n}$ of the algorithm, but also in the dual $Y^{n}$, as in (2.4). Unfortunately, the next theorem shows that the time-continuous limit (2.8) always remains asymptotically inconsistent.
Theorem 3.6.
Consider the same target $\mu^{\ast}$ as in Theorem 3.5, but the more general diffusion (2.8)
$\mathrm{d}Z_{t}=-A_{\text{{PD}}}Z_{t}\mathrm{d}t+B_{\text{{gPD}}}\mathrm{d}W_{t},\quad\text{with }A_{\text{{PD}}}=\begin{pmatrix}c_{g}^{-1}&k\\ -\lambda k&\lambda c_{f}\end{pmatrix},B_{\text{{gPD}}}=\begin{pmatrix}b_{11}&b_{12}\\ b_{21}&b_{22}\end{pmatrix}$ (3.9)
where $W_{t}$ is Brownian motion in $\mathbb{R}^{2}$. Then there exists no choice for the entries of $B_{\text{{gPD}}}\in\mathbb{R}^{2\times 2}$ independent of $c_{g},c_{f}$ and $k$ such that the stationary solution $\pi_{\lambda}$ of (3.9) has $x$-marginal $\mu^{\ast}$.
Proof.
We start by noting that the contraction result Theorem 3.4 can be applied also the more general $\Sigma_{\text{{gPD}}}=\frac{1}{2}B_{\text{{gPD}}}B_{\text{{gPD}}}^{T}$ with any constant coefficient $B_{\text{{gPD}}}$, since the relevant Lemma 3.3 holds for any constant diffusion coefficient. In view of this contraction property, we can assume the initialization $Z_{0}\sim\pi_{0}$ to be Gaussian with mean $z_{0}$, covariance $\Sigma_{0}$ and obtain $\pi_{\lambda}$ by computing $\lim_{t\to\infty}\pi_{t}$. The reason for initializing at a Gaussian is that we can again obtain the closed-form solution $\pi_{t}$ of the Cauchy problem for the Fokker-Planck equation. We carry out the same computations as in the proof of Theorem 3.5, showing that $\pi_{t}$ is Gaussian with mean $E(t)^{T}z_{0}$ and covariance matrix
$E(t)^{T}\Sigma_{0}E(t)+2\int_{0}^{t}E(s)^{T}B_{\text{{gPD}}}B_{\text{{gPD}}}^{T}E(s)\,\mathrm{d}s.$
The entries of $E$ can be computed explicitly (we omit the terms here) and by integrating and taking the limit $t\to\infty$, we find that the $x$-marginal of the stationary distribution $\pi_{\lambda}$ has variance
$\operatorname{Var}[\mu_{\lambda}]=\frac{(b_{11}^{2}+b_{12}^{2})(\lambda c_{f}^{2}c_{g}^{2}+c_{g}(c_{f}+k^{2}c_{g}))-2(b_{11}b_{21}+b_{12}b_{22})kc_{f}c_{g}^{2}+(b_{21}^{2}+b_{22}^{2})\lambda^{-1}k^{2}c_{g}^{2}}{2(1+\lambda c_{f}c_{g})(c_{f}+k^{2}c_{g})}.$
We want to show that there is no choice of $b_{ij}$ (independent of $c_{f},c_{g}$ and $k$) such that the target $\mu^{\ast}$ coincides with the $x$-marginal $\mu_{\lambda}$ of the stationary solution. The variance of both distributions is equal if and only if
$2c_{g}c_{f}(1+\lambda c_{g}c_{f})=(b_{11}^{2}+b_{12}^{2})(\lambda c_{f}^{2}c_{g}^{2}+c_{g}(c_{f}+k^{2}c_{g}))-2(b_{11}b_{21}+b_{12}b_{22})kc_{f}c_{g}^{2}+(b_{21}^{2}+b_{22}^{2})\lambda^{-1}k^{2}c_{g}^{2}.$
Isolating the terms containing $k^{2}$ shows that this implies
$0=(b_{11}^{2}+b_{12}^{2})c_{g}^{2}+(b_{21}^{2}+b_{22}^{2})\lambda^{-1}c_{g}^{2},$
which is only satisfied if all $b_{ij}$ are zero since $\lambda>0$. Hence, a problem-independent choice for $B_{\text{{cPD}}}$ that results in the target being invariant with respect to the diffusion does not exist. ∎
We want to elaborate more on the concentration properties of $\pi_{\lambda}$ in the scheme (2.6) with primal diffusion. Theorem 3.5 showed that in the limiting case $\lambda\to\infty$, the stationary distribution does not have a Lebesgue density anymore since the mass concentrates on the optimality condition $\partial f^{\ast}(y)=Kx$. In that case, the primal marginal actually converges to the target distribution, which is intuitively clear since inserting $Y_{t}\in\partial f(KX_{t})$ into the equation for $X_{t}$ reduces the system to overdamped Langevin diffusion. We will rigorously generalize these properties of the toy example of Theorem 3.5 in two ways. Specifically, we show in Section 3.3 under which conditions the stationary solution converges, as $\lambda\to\infty$, to a singular push-forward of the target $\mu^{\ast}$ in the space of distributions over $\mathcal{Z}$. Also, we next show that the concentration at the optimal dual value generally does not occur for finite $\lambda$, even though the diffusion coefficient is degenerate.
In the case of smooth coefficients, the Fokker-Planck operator can be shown to be hypoelliptic, implying that any solution to the dynamics has a smooth density with respect to the Lebesgue measure for all times $t>0$, even in the case of a concentrated initialization.
Definition 3.7.
Let $\mathcal{L}$ be a partial differential operator with smooth coefficients. If, for every open set $\Omega$ and every distribution $u$ on $\Omega$, $\mathcal{L}u\in C^{\infty}(\Omega)$ implies $u\in C^{\infty}(\Omega)$, then $\mathcal{L}$ is called hypoelliptic.
Every elliptic operator is hypoelliptic, and equally any operator of the form $-\partial_{t}+\mathcal{L}$ where $\mathcal{L}$ is the generator of a diffusion process with uniformly positive definite diffusion coefficient. A special case is overdamped Langevin diffusion with smooth drift, which naturally has a smooth transition density.
The degenerate diffusion coefficient in (2.6) results in a constant but only positive semidefinite diffusion coefficient $\Sigma_{\text{{PD}}}$ in the Fokker-Planck equation (3.8). In order to prove that the Fokker-Planck equation is hypoelliptic, we hence have to employ a weaker condition for hypoellipticity, usually termed ‘Hörmander’s condition’ [Hoermander1967]. In order to state it, we need the definition of the commutator operation and Lie algebras spanned by smooth vector fields.
Definition 3.8.
For two vector fields $A,B$ on $\mathbb{R}^{n}$, the Lie bracket or commutator $[A,B]$ is the vector field on $\mathbb{R}^{n}$ defined by
$[A,B](x)=DB(x)A(x)-DA(x)B(x),$
with the derivative matrices $(DA)_{i,j}(x)=\partial_{j}A_{i}(x)$.
Lemma 3.9 ([Hoermander1967]).
Assume that the generator $\mathcal{L}$ of a diffusion process with state space $\mathbb{R}^{n}$ has $C^{\infty}$-smooth coefficients and can be written in the form
$\mathcal{L}=A_{0}+\sum_{j=1}^{n}A_{j}^{2}$
with first-order differential operators $A_{0},\dots,A_{n}$. Interpreting the differential operators as vector fields, define a collection of vector fields $\mathcal{A}_{k}$ by
$\mathcal{A}_{0}\coloneqq\{A_{j}\,:\,j=1,\dots,n\},\quad\mathcal{A}_{k+1}=\mathcal{A}_{k}\cup\{[A,A_{j}]\,:\,A\in\mathcal{A}_{k},j=0,\dots,n\}.$
If the vector spaces $\mathcal{A}_{k}(x)\coloneqq\operatorname{span}\left(\{A(x)\,:\,A\in\mathcal{A}_{k}\}\right)$ satisfy $\bigcup_{k\geq 0}\mathcal{A}_{k}(x)=\mathbb{R}^{n}$ for every $x\in\mathbb{R}^{n}$, then $\mathcal{L}$ is said to satisfy the parabolic Hörmander condition. In that case, $\partial_{t}-\mathcal{L}$ and $\partial_{t}-\mathcal{L}^{\ast}$ are both hypoelliptic, hence any process generated by $\cal L$ has $C^{\infty}$-smooth transition probability. In particular, any solution to the Fokker-Planck equation $\partial_{t}p=\mathcal{L}^{\ast}p$ is $C^{\infty}((0,\infty)\times\mathbb{R}^{n})$-smooth.
We refer to Chapter 6 of the book [Pavliotis2014] for a discussion in the context of diffusion SDEs and an application to the Langevin diffusion SDE with friction. For the diffusion equation arising in our primal-dual setting, the generator is given by
$\mathcal{L}_{\text{{PD}}}u=-A_{\text{{PD}}}^{T}(z)\nabla_{z}u+\Delta_{x}u.$
The following result gives a sufficient condition under which the primal-dual dynamics considered here satisfy Hörmander’s condition, so that the corresponding Fokker-Planck operator is hypoelliptic.
Theorem 3.10.
Assume that $g\in C^{\infty}(\mathcal{X})$ and $f^{\ast}\in C^{\infty}(\mathcal{Y})$, so that the diffusion (2.6) has smooth coefficients. If $\dim(\mathcal{Y})=m\leq d=\dim(\mathcal{X})$ and $K$ has full rank $m$, then $\mathcal{L}_{\text{{PD}}}$ satisfies Hörmander’s condition. As a consequence, the stationary solution $\pi_{\lambda}$ has a $C^{\infty}(\mathcal{Z})$-smooth Lebesgue density.
Proof.
We apply Lemma 3.9. $\mathcal{L}_{\text{{PD}}}$ has the correct form with vector fields $A_{j}=\partial_{x_{j}}$, $j=1,\dots,d$ and
$A_{0}=-A_{\text{{PD}}}^{T}(z)\nabla_{z}=\sum_{j=1}^{d}\left(-\frac{\partial g}{\partial x_{j}}(x)-k_{j}^{T}y\right)\partial_{x_{j}}+\lambda\sum_{i=1}^{m}\left((k^{i})^{T}x-\frac{\partial f^{\ast}}{\partial y_{i}}(y)\right)\partial_{y_{i}},$
where we denoted the columns of $K$ by $k_{1},\dots,k_{d}$ and the rows by $k^{1},\dots,k^{m}$. Clearly, the vector fields $A_{1},\dots,A_{k}$ span the first $d$ dimensions $\partial_{x_{j}}$, $j=1,\dots,d$, i.e. $\mathcal{A}_{0}(z)=\mathcal{X}$ for all $z\in\mathcal{Z}$. The Lie brackets in $\mathcal{A}_{1}$ are given by
$[A_{j},A_{0}]=-(H_{g}(x))_{j}^{T}\nabla_{x}-\lambda k_{j}^{T}\nabla_{y},$
where $H_{g}$ is the Hessian of $g$. Due to $\lambda>0$ and the full rank assumption, the columns $k_{j}$ of $K$ span $\mathcal{Y}$ and we have $\mathcal{A}_{1}(z)=\mathcal{Z}$ for all $z\in\mathcal{Z}$. ∎
While the smoothness assumption on $g,f^{\ast}$ is necessary for technical reasons, our numerical test validated that the dispersion in the dual variable is a more general phenomenon. Concentration in the dual variable can generally not to be expected in the primal-dual dynamics.
3.3 Convergence to Overdamped Langevin Diffusion
We now continue by showing that in the limiting case $\lambda\to\infty$ and under some smoothness assumption on $f$, the stationary solution $\pi_{\lambda}$ of (2.6) becomes singular. In the limit, the primal marginal distribution coincides with the target.
Consider a solution $(X_{t},Y_{t})$ of (2.6) with distribution $(X_{t},Y_{t})\sim\pi_{t}$, denoting the $x$-marginal by $X_{t}\sim\mu_{t}$. We are interested in the behaviour of $\mu_{t}$ in the limit $\lambda\rightarrow\infty$. Intuitively, as $\lambda$ becomes large, it is clear that $Y_{t}$ is forced to concentrate, with the (formally) limiting SDE taking the form
$dX_{t}$ $\in-(\partial g(X_{t})+K^{T}Y_{t}dt+\sqrt{2}dW_{t}$
$Y_{t}$ $\in\partial f(KX_{t})).$
This coincides with overdamped Langevin diffusion on the potential $g+f\circ K$, with the target $\mu^{\ast}$ as invariant distribution.
This suggests we can use a coupling argument between (3.3) and the primal-dual dynamics (2.6) to bound the Wasserstein distance between the stationary state of (2.6) and the target.
Theorem 3.11.
Assume that $f\in C^{3}$, that $g,f^{\ast}$ are $\omega_{g}$- and $\omega_{f^{\ast}}$-strongly convex, respectively, and that the target distribution $\mu^{\ast}$ satisfies
$\mathbb{E}_{X\sim\mu^{\ast}}\left[\lVert\nabla(\nabla f\circ K)(X)\rVert_{F}^{2}\right]=:D_{1}<\infty,$
$\mathbb{E}_{X\sim\mu^{\ast}}\left[\lVert\nabla\cdot\nabla(\nabla f\circ K)(X)\rVert^{2}\right]=:D_{2}<\infty,$
$\mathbb{E}_{X\sim\mu^{\ast}}\left[\lVert\nabla(\nabla f\circ K)(X)\left(\nabla g(X)+K^{T}\nabla f(KX)\right)\rVert^{2}\right]=:D_{3}<\infty,$
with constants $D_{i}$, where in the second line the Laplacian is understood acting component-wise on the $\mathcal{Y}$-valued function $\nabla f\circ K$. For $\omega_{g}/\omega_{f^{\ast}}<\lambda$ and a constant $C_{1}(\lambda)=(\lambda\omega_{g}/D_{1})^{-1/2}+\mathcal{O}(\lambda^{-3/2})$ as $\lambda\to\infty$, the following two statements hold:
  1. 1.
    If the primal-dual dynamics (2.6) are initialized with $X_{0}\sim\mu_{0}$ for some $\mu_{0}\in\mathcal{P}_{2}(\mathcal{X})$ and $Y_{0}$ concentrated at the optimality condition $Y_{0}=\nabla f(KX_{0})$, then for every $t>0$ it holds
    $\mathcal{W}_{2}(\mu_{t},\mu^{\ast})\leq\left((1+\lambda^{-1}\omega_{f^{\ast}}\lVert K\rVert)^{1/2}\mathcal{W}_{2}(\mu_{0},\mu^{\ast})+C_{1}(\lambda)\right)e^{-\omega_{g}t}+C_{1}(\lambda).$
  2. 2.
    The $x$-marginal $\mu_{\lambda}$ of the stationary solution $\pi_{\lambda}$ of (2.6) satisfies
    $\mathcal{W}_{2}(\mu_{\lambda},\mu^{\ast})\leq C_{1}(\lambda).$
Proof.
Assume $(X_{t},Y_{t})$ with distribution $\pi_{t}$ solves (2.5). Now consider a coupling with a process $\tilde{X}_{t}$ which is initialized as $\tilde{X}_{0}\sim\mu^{\ast}$ and solves (3.3) with the same Brownian motion $W_{t}$ as in (2.5). Since $\mu^{\ast}$ is the invariant distribution of (3.3), it follows $\tilde{X}_{t}\sim\mu^{\ast}$ for all $t>0$. The joint measure $\Pi_{t}$ of $(X_{t},Y_{t},\tilde{X}_{t})$ solves the Fokker-Planck equation
$\frac{\partial}{\partial t}\Pi_{t}$ $=\nabla_{x}\cdot((p+K^{T}y)\Pi_{t})+\lambda\nabla_{y}\cdot((q-Kx)\Pi_{t})+\nabla_{\tilde{x}}\cdot((\tilde{p}+K^{T}\tilde{y})\Pi_{t})$ (3.10)
$\qquad+(\nabla_{x}+\nabla_{\tilde{x}})\cdot(\nabla_{x}+\nabla_{\tilde{x}})\Pi_{t}.$
where as before $p\in\partial g(x),q\in\partial f^{\ast}(y),\tilde{p}\in\partial g(\tilde{x})$. With the artificial, concentrated dual variable $\tilde{Y}_{t}=\nabla f(K\tilde{X}_{t})$, we denote the singular joint law of $(\tilde{X}_{t},\tilde{Y}_{t})$ by $\pi^{\ast}=(I,\nabla f\circ K)_{\#}\mu^{\ast}$.
We now consider the transport distance $\mathcal{T}_{1,\lambda^{-1}}$ on $\mathcal{P}(\mathcal{Z})$ in order to bound the distance between $\pi_{t}$ and $\pi^{\ast}$, which due to the concentration of $\pi^{\ast}$ can be written as
$\mathcal{T}_{1,\lambda^{-1}}(\pi_{t},\pi^{\ast})=\inf_{\Pi\in\Gamma(\pi,\mu^{\ast})}\left(\mathbb{E}_{(x,y,\tilde{x})\sim\Pi}\left[d_{1,\lambda^{-1}}(x,y,\tilde{x})^{2}\right]\right)^{1/2},$
with an augmented distance function
$d_{1,\lambda^{-1}}(x,y,\tilde{x}):=\left(\lVert x-\tilde{x}\rVert^{2}+\lVert y-\nabla f(K\tilde{x})\rVert_{\lambda^{-1}}^{2}\right)^{1/2}.$
By using the joint Fokker-Planck equation (3.10) and the notations $\tilde{y}(\tilde{x})\coloneqq\nabla f(K\tilde{x})$ as well as $M\coloneqq\nabla\tilde{y}$ and $m\coloneqq\nabla\cdot\nabla\tilde{y}$, we obtain
$\frac{1}{2}\frac{\,\mathrm{d}}{\,\mathrm{d}t}\mathbb{E}_{(x,y,\tilde{x})\sim\Pi_{t}}\left[d_{1,\lambda^{-1}}(x,y,\tilde{x})^{2}\right]$
$=\frac{1}{2}\int d_{1,\lambda^{-1}}(x,y,\tilde{x})^{2}\left(\nabla_{x}\cdot((p+K^{T}y)\Pi_{t})+\lambda\nabla_{y}\cdot((q-Kx)\Pi_{t})+\nabla_{\tilde{x}}\cdot((\tilde{p}+K^{T}\tilde{y}(\tilde{x}))\Pi_{t})\right)\mathrm{d}(x,y,\tilde{x})$
$\qquad+\frac{1}{2}\int d_{1,\lambda^{-1}}(x,y,\tilde{x})^{2}\left((\nabla_{x}+\nabla_{\tilde{x}})\cdot(\nabla_{x}+\nabla_{\tilde{x}})\Pi_{t}\right)\,\mathrm{d}(x,y,\tilde{x})$
$=-\int\left((x-\tilde{x})\cdot(p+K^{T}y)+(y-\tilde{y}(\tilde{x}))\cdot(q-Kx)+(\tilde{x}-x)\cdot(\tilde{p}+K^{T}\tilde{y}(\tilde{x})\right)\,\mathrm{d}\Pi_{t}$
$\qquad-\frac{1}{\lambda}\int\left(M(\tilde{x})(\tilde{y}(\tilde{x})-y)\right)\cdot(\tilde{p}+K^{T}\tilde{y}(\tilde{x}))\,\mathrm{d}\Pi_{t}-\frac{1}{\lambda}\int\left(M(\tilde{x})(\tilde{y}(\tilde{x})-y)\right)\cdot(\nabla_{x}+\nabla_{\tilde{x}})\Pi_{t}d(x,y,\tilde{x})$
$=-\int(x-\tilde{x})\cdot(p-\tilde{p})\,\mathrm{d}\Pi_{t}-\int(y-\tilde{y}(\tilde{x}))\cdot(q-K\tilde{x})\,\mathrm{d}\Pi_{t}$
$\qquad+\frac{1}{\lambda}\int\left[(\tilde{y}(\tilde{x})-y)\cdot\left(m(\tilde{x})-M^{T}(\tilde{x})(\tilde{p}+K^{T}\tilde{y}(\tilde{x}))\right)+\lVert M(\tilde{x})\rVert_{F}^{2}\right]\,\mathrm{d}\Pi_{t}$
For the terms in the second to last line, we can use strong convexity of $g$ and $f^{\ast}$ to estimate
$\int(x-\tilde{x})\cdot(p-\tilde{p})\,\mathrm{d}\Pi_{t}\geq\omega_{g}\mathbb{E}_{(x,y,\tilde{x})\sim\Pi_{t}}\left[\lVert x-\tilde{x}\rVert^{2}\right]$
$\int(y-\tilde{y}(\tilde{x}))\cdot(q-K\tilde{x})\,\mathrm{d}\Pi_{t}\geq\lambda\omega_{f^{\ast}}\mathbb{E}_{(x,y,\tilde{x})\sim\Pi_{t}}\left[\lVert y-\tilde{y}(\tilde{x})\rVert_{\lambda^{-1}}^{2}\right]$
Since the $\tilde{x}$-marginal of $\Pi_{t}$ is the target $\mu^{\ast}$, by assumption we have
$\lambda^{-1}\int\lVert M(\tilde{x})\rVert_{F}^{2}\,\mathrm{d}\Pi_{t}\leq\lambda^{-1}D_{1}.$
In a similar way, using the Cauchy-Schwarz inequality and Young’s inequality, for any $\delta>0$ we can bound the final term by
$\lambda^{-1}\int(\tilde{y}(\tilde{x})-y)$ $\cdot\left(m(\tilde{x})-M^{T}(\tilde{x})(\tilde{p}+K^{T}\tilde{y}(\tilde{x}))\right)\,\mathrm{d}\Pi_{t}$
$\leq\lambda^{-1}\int\lVert\tilde{y}(\tilde{x})-y\rVert\cdot\lVert m(\tilde{x})-M^{T}(\tilde{x})(\tilde{p}+K^{T}\tilde{y}(\tilde{x}))\rVert\,\mathrm{d}\Pi_{t}$
$\leq\frac{\delta}{2}\mathbb{E}_{(x,y,\tilde{x})\sim\Pi_{t}}\left[\lVert y-\tilde{y}(\tilde{x})\rVert_{\lambda^{-1}}^{2}\right]+\frac{1}{2\delta\lambda}\mathbb{E}_{\tilde{x}\sim\mu^{\ast}}\left[\lVert m(\tilde{x})-M^{T}(\tilde{x})(\tilde{p}+K^{T}\tilde{y}(\tilde{x}))\rVert^{2}\right]$
$\leq\frac{\delta}{2}\mathbb{E}_{(x,y,\tilde{x})\sim\Pi_{t}}\left[\lVert y-\tilde{y}(\tilde{x})\rVert_{\lambda^{-1}}^{2}\right]+\frac{D_{2}+D_{3}}{2\delta\lambda}.$
Denoting $\varphi(t):=\mathbb{E}_{(x,y,\tilde{x})\sim\Pi_{t}}\left[d_{1,\lambda^{-1}}(x,y,\tilde{x})^{2}\right]$, we therefore have
$\varphi^{\prime}(t)\leq-2\omega\varphi(t)+\lambda^{-1}(2D_{1}+(D_{2}+D_{3})/\delta),$
where we denoted $\omega\coloneqq\omega(\delta)\coloneqq\min\{\omega_{g},\lambda\omega_{f^{\ast}}-\frac{\delta}{2}\}$. Using Gronwall’s lemma, this implies
$\varphi(t)$ $\leq e^{-2\omega t}\varphi(0)+\int_{0}^{t}e^{-2\omega(t-s)}\lambda^{-1}(2D_{1}+(D_{2}+D_{3})/\delta)\,\mathrm{d}s$
$=e^{-2\omega t}\varphi(0)+\lambda^{-1}(2D_{1}+(D_{2}+D_{3})/\delta)\frac{1-e^{-2\omega t}}{2\omega}$
By the definition of $\mathcal{T}_{1,\lambda^{-1}}$ and taking the infimum over all possible initializations $\Pi_{0}$ with $(x,y)$-marginal $\pi_{0}$ and $(\tilde{x},\tilde{y})$-marginal $\pi^{\ast}=(id,\nabla f\circ K)_{\#}\mu^{\ast}$ on the right side, we get
$\mathcal{T}_{1,\lambda^{-1}}^{2}(\pi_{t},\pi^{\ast})\leq\varphi(t)\leq e^{-2\omega t}\mathcal{T}_{1,\lambda^{-1}}^{2}(\pi_{0},\pi^{\ast})+\lambda^{-1}(2D_{1}+(D_{2}+D_{3})/\delta)\frac{1-e^{-2\omega t}}{2\omega}.$ (3.11)
Bearing in mind the dependency of $\omega$ on $\delta$, for $\lambda<\omega_{g}/\omega_{f^{\ast}}$ the bound on the right is minimized by $\delta=2(\lambda\omega_{f^{\ast}}-\omega_{g})$ for any finite $t$ as well as in the limit $t\to\infty$. This choice implies $\omega=\omega_{g}$ and we denote the resulting constant terms by
$C(\lambda)\coloneqq\left(\frac{D_{1}}{\lambda\omega_{g}}+\frac{D_{2}+D_{3}}{4\omega_{g}(\lambda\omega_{f^{\ast}}-\omega_{g})}\right)^{1/2}=(\lambda\omega_{g}/D_{1})^{-1/2}+\mathcal{O}(\lambda^{-3/2}).$
The first desired inequality for $t>0$ follows by taking the square root of (3.11) and using that $\mathcal{W}_{2}(\mu_{t},\mu^{\ast})\leq\mathcal{T}_{1,\lambda^{-1}}(\pi_{t},\pi^{\ast})$ and, assuming a concentrated initialization $Y_{0}=\nabla f(KX_{0})$, that
$\mathcal{T}_{1,\lambda^{-1}}(\pi_{0},\pi^{\ast})\leq(1+\lambda^{-1}\omega_{f^{\ast}}\lVert K\rVert)^{1/2}\mathcal{W}_{2}(\mu_{0},\mu^{\ast}).$
Since $\pi_{t}$ converges to the stationary distribution $\pi_{\lambda}$, taking the limit $t\to\infty$ in (3.11) and using $\mathcal{W}_{2}(\mu_{\lambda},\mu^{\ast})\leq\mathcal{T}_{1,\lambda^{-1}}(\pi_{\lambda},\pi^{\ast})$ proves the second statement. ∎
3.4 A Bias-Corrected Modification of Primal-Dual Diffusion
In this section, we want to follow up on the negative result of Theorem 3.6, where we saw that there is no way to construct a diffusion SDE with the primal-dual type drift coefficient and homogeneous diffusion that has the correct stationary distribution. Naturally, we can also consider modifications of the diffusion SDE with problem-dependent and/or non-homogeneous diffusion coefficient, so that the samples become (in continuous time) unbiased.
We aim to correct the Fokker-Planck equation (3.8) by adding terms such that $\pi^{\ast}=(I,\nabla f\circ K)_{\#}\mu^{\ast}$ becomes the stationary solution of the PDE. Throughout this subsection, we will assume that $g\in C^{1}(\mathbb{R}^{d})$ and $f\in C^{3}(\mathbb{R}^{m})$, we will denote as before $M:\mathbb{R}^{d}\to\mathbb{R}^{d\times m}$ for $M(x)\coloneqq\nabla(\nabla f\circ K)^{T}(x)=K^{T}H_{f}(Kx)$. Instead of (2.5), consider the modified stochastic differential equation
$\,\mathrm{d}Z_{t}=-A_{\text{{mod}}}(Z_{t})\,\mathrm{d}t+B_{\text{{mod}}}(Z_{t})\,\mathrm{d}W_{t},$ (3.12)
where the drift and diffusion coefficients are given by
$A_{\text{{mod}}}(z)$ $\coloneqq\begin{pmatrix}\nabla g(x)+K^{T}y\\ \lambda(\nabla f^{*}(y)-Kx)+M(x)^{T}(\nabla g(x)+K^{T}\nabla f(Kx))\end{pmatrix}$
$B_{\text{{mod}}}(z)$ $\coloneqq\sqrt{2}\,\nabla(id,\nabla f\circ K)(x)=\sqrt{2}\begin{pmatrix}I\\ M(x)^{T}\end{pmatrix}.$
Note that $B_{\text{{mod}}}$ depends on $x$ and is therefore not a special case of $B_{\text{{gPD}}}$. The drift $A_{\text{{mod}}}$ contains the same terms as $A_{\text{{PD}}}$, and an additional term $-M(x)^{T}\nabla\log\mu^{\ast}(x)$.
Remark 3.12.
As an intuitive explanation why (3.12) has the desired stationary distribution, note that we want the stationary solution to concentrate its mass in $\mathcal{Z}=\mathbb{R}^{d+m}$ on the $d$-dimensional manifold $\mathcal{U}$ described by $y=\nabla f(Kx)$. The tangent space to $\mathcal{U}$ at any point $(x,\nabla f(Kx))$, denoted in the following by $\mathcal{T}_{x}\mathcal{U}$, is spanned by the columns of $\nabla(id,\nabla f\circ K)=(I,M(x))^{T}$. This coincides with the column space of the modified diffusion coefficient $\Sigma_{\text{{mod}}}\coloneqq\frac{1}{2}B_{\text{{mod}}}B_{\text{{mod}}}^{T}$, hence locally on $\mathcal{U}$, the diffusion is confined to the manifold. The term $\nabla f^{\ast}(y)-Kx$ in the drift vanishes upon concentration of the mass on $\mathcal{U}$, the remaining terms are just the gradient of the negative log-density $-\nabla\log\mu^{\ast}(x)=\nabla g(x)+K^{T}\nabla f(Kx)$ in the primal variable $x$ mapped to the tangent space of $\mathcal{U}$ by the map $(I,M(x))^{T}$. The drift hence does not transport mass away from $\mathcal{U}$ and locally on $\mathcal{U}$, (3.12) coincides with overdamped Langevin diffusion.
In order to show formally that the partially singular distribution $\pi^{\ast}$ is a stationary solution of (3.12), we have to consider the weak formulation of the corresponding Fokker-Planck equation, given by
[…]
Theorem 3.13.
The distribution $\pi^{\ast}\coloneqq(I,\nabla f\circ K)_{\#}\mu^{*}$ is a stationary solution of (3.13).
In principle, this theorem might suggest that a time discretization of the modified (3.12) may be a better idea than the primal-dual sampling algorithm (1.4). Note, however, apart from the higher-order derivative terms in $f$ (which counter the motivation to use a primal-dual scheme in the first place), the term $A_{\text{{mod}}}$ contains just the gradient of the target log-density, i.e. we explicitly add overdamped Langevin diffusion in the dual variable. This suggests that, algorithmically, the dualization is somewhat redundant and any sampler based on (3.12) would be inferior to a standard unadjusted Langevin algorithm.
In order to prove Theorem 3.13, we need the following vector calculus identity.
Lemma 3.14.
Let $\psi:\mathbb{R}^{d}\times\mathbb{R}^{m}\rightarrow\mathbb{R}^{n}$, $\eta:\mathbb{R}^{d}\rightarrow\mathbb{R}^{m}$ and $S:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d\times n}$, each with $C^{1}$ entries. Define $r:\mathbb{R}^{d}\to\mathbb{R}^{n}$ by $r_{j}(x)\coloneqq\sum_{i}\frac{\partial S_{ij}}{\partial x_{i}}(x)$ and denote the partial Jacobians of $\psi$ by $D_{x}\psi:\mathbb{R}^{d}\times\mathbb{R}^{m}\to\mathbb{R}^{n\times d}$ with $(D_{x}\psi)_{i,j}(x,y)\coloneqq\frac{\partial\psi_{i}}{\partial x_{j}}(x,y)$ and $D_{y}\psi:\mathbb{R}^{d}\times\mathbb{R}^{m}\to\mathbb{R}^{n\times m}$ where $(D_{y}\psi)_{i,k}(x,y)\coloneqq\frac{\partial\psi_{i}}{\partial y_{k}}(x,y)$. Then for all $x\in\mathbb{R}^{d}$ it holds
$\nabla_{x}\cdot(S(x)\psi(x,\eta(x)))=r(x)\cdot\psi(x,\eta(x))+\operatorname{tr}(S(x)D_{x}\psi(x,\eta(x))+\operatorname{tr}(S(x)D_{y}\psi(x,\eta(x))D_{x}\eta(x)).$ (3.14)
Proof.
The proof is a straightforward application of product and chain rule.
$\nabla_{x}\cdot($ $S(x)\psi(x,\eta(x)))$
$=\sum_{i=1}^{d}\frac{\partial}{\partial x_{i}}(\sum_{j=1}^{d}S_{ij}(x)\psi_{j}(x,\eta(x)))$
$=\sum_{ij}\frac{\partial S_{ij}}{\partial x_{i}}(x)\psi_{j}(x,\eta(x))+S_{ij}(x)\frac{\partial\psi_{j}}{\partial x_{i}}(x,\eta(x))+S_{ij}(x)\sum_{k=1}^{m}\frac{\partial\psi_{j}}{\partial y_{k}}(x,\eta(x))\frac{\partial h_{k}}{\partial x_{i}}(x)$
$=r(x)\cdot\psi(x,\eta(x))+\operatorname{tr}(S(x)D_{x}\psi(x,\eta(x))+\operatorname{tr}(S(x)D_{y}\psi(x,\eta(x))D_{x}\eta(x)).$
Proof of Theorem 3.13.
It suffices to check that the right side of (3.13) evaluates to zero for any $\varphi\in C^{\infty}(\mathbb{R}^{d}\times\mathbb{R}^{d})$ when we insert $\pi^{\ast}$:
$\int\begin{pmatrix}\nabla_{x}\\ \nabla_{y}\end{pmatrix}^{T}\begin{pmatrix}I&M(x)\\ M(x)^{T}&M(x)M(x)^{T}\end{pmatrix}\begin{pmatrix}\nabla_{x}\\ \nabla_{y}\end{pmatrix}\varphi\,\mathrm{d}\pi^{\ast}-\int\nabla_{x}\varphi\cdot(\nabla g(x)+K^{T}y)\,\mathrm{d}\pi^{\ast}$ (3.15a)
$-\int\nabla_{y}\varphi\cdot(\lambda\nabla f^{*}(y)-\lambda Kx)\,\mathrm{d}\pi^{\ast}-\int\nabla_{y}\varphi\cdot M(x)^{T}(\nabla g(x)+K^{T}\nabla f(Kx))\,\mathrm{d}\pi^{\ast}$ (3.15b)
Using $(\nabla f)^{-1}=\nabla f^{\ast}$ and the fact that $\pi^{\ast}$ is partially concentrated on $y=\nabla f(Kx)$, it follows that
$\int\nabla_{y}\varphi\cdot(\lambda\nabla f^{\ast}(y)-\lambda Kx)\,\mathrm{d}\pi^{\ast}=0.$
Using partial integration and Lemma 3.14 with $S\equiv I$, $\psi=\nabla_{x}\varphi$ and $\eta=\nabla f\circ K$, the second integral in (3.15a) can be rewritten as
$-\int\nabla_{x}$ $\varphi\cdot(\nabla g(x)+K^{T}y)\,\mathrm{d}\pi^{\ast}$
$=\int\nabla_{x}\varphi(x,\nabla f(Kx))\cdot(-\nabla g(x)-K^{T}\nabla f(Kx))\exp(-f(Kx)-g(x))\,\mathrm{d}x$
$=\int\nabla_{x}\varphi(x,\nabla f(Kx))\cdot\nabla_{x}(\exp(-f(Kx)-g(x)))\,\mathrm{d}x$
$=-\int\nabla_{x}\cdot(\nabla_{x}\varphi(x,\nabla f(Kx)))\,\mathrm{d}\mu^{\ast}(x)$
$=-\int\operatorname{tr}(D_{x}\nabla_{x}\varphi(x,\nabla f(Kx)))\,\mathrm{d}\mu^{\ast}(x)-\int\operatorname{tr}(D_{y}\nabla_{x}\varphi(x,\nabla f(Kx))M(x)^{T})\,\mathrm{d}\mu^{\ast}(x)$
$=-\int\Delta_{x}\varphi(x,\nabla f(Kx))\,\mathrm{d}\mu^{\ast}(x)-\int\nabla_{y}\cdot(M(x)^{T}\nabla_{x}\varphi(x,\nabla f(Kx)))\,\mathrm{d}\mu^{\ast}(x)$
where we used that $M(x)^{T}=D_{x}\eta(x)$ for $\eta=\nabla f\circ K$. These two terms cancel with the terms from the first column of the diffusion matrix in (3.15a). The second integral in (3.15b) can also be reformulated using partial integration and then Lemma 3.14, this time with $S=M$, $\psi=\nabla_{y}\varphi$ and $\eta=\nabla f\circ K$. In the notation of Lemma 3.14, we get $r:\mathbb{R}^{d}\to\mathbb{R}^{m}$ with $r_{j}(x)\coloneqq\sum_{i}\frac{\partial M_{ij}}{\partial x_{i}}(x)$ and obtain
-$\int$- ∫ $\nabla_{y}\varphi(x,y)\cdot M(x)^{T}(\nabla g(x)+K^{T}\nabla f(Kx))\,\mathrm{d}\pi^{\ast}$
$=\int\nabla_{y}\varphi(x,\nabla f(Kx))\cdot M(x)^{T}(-\nabla g(x)-K^{T}\nabla f(Kx))\exp(-f(Kx)-g(x))\,\mathrm{d}x$
$=\int\nabla_{y}\varphi(x,\nabla f(Kx))\cdot M(x)^{T}\nabla_{x}(\exp(-f(Kx)-g(x)))\,\mathrm{d}x$
$=-\int\nabla_{x}\cdot(M(x)\nabla_{y}\varphi(x,\nabla f(Kx)))\,\mathrm{d}\mu^{\ast}(x)$
$=-\int r(x)\cdot\nabla_{y}\varphi(x,\nabla f(Kx))\,\mathrm{d}\mu^{\ast}(x)-\int\operatorname{tr}(M(x)D_{x}\nabla_{y}\varphi(x,\nabla f(Kx)))\,\mathrm{d}\mu^{\ast}(x)$
$\qquad\qquad-\int\operatorname{tr}(M(x)D_{y}\nabla_{y}\varphi(x,\nabla f(Kx))M(x)^{T})\,\mathrm{d}\mu^{\ast}(x)$
Since $\operatorname{tr}(MD_{y}\nabla_{y}\varphi M^{T})=\operatorname{tr}(M^{T}MD_{y}\nabla_{y}\varphi)=\nabla_{y}\cdot(M^{T}M\nabla_{y}\varphi)$, the term in the last line cancels with the term from the lower right entry in the diffusion matrix. By the product rule, $\nabla_{x}\cdot(M\nabla_{y}\varphi)=\operatorname{tr}(MD_{x}\nabla_{y}\varphi)+r\cdot\nabla_{y}\varphi$, hence the second to last line cancels with the term from the upper right entry in the diffusion matrix. It follows that (3.15) is equal to zero. ∎
Remark 3.15.
While the the last result showed that the target is now the invariant solution, the asymptotic stability of the new diffusion process is unclear. In general, it is hard to generalize a contraction result like Lemma 3.3 to the case of non-homogeneous diffusion coefficients. In section 2 of [Fournier2021], this question is addressed in more detail for the heat equation with non-homogeneous diffusion. The authors show that it becomes possible to prove stability in a Wasserstein distance with a modified underlying metric, where the metric depends on the diffusion coefficient via an elliptic PDE.
Since the modified equation (3.12) is of limited numerical interest, we only sketch some main ideas and omit a detailed proof. At every point $x\in\mathcal{X}$ one can define an inner product $\langle u,v\rangle_{x}\coloneqq u^{T}V^{-1}(x)v$ where $V(x)=W(x)W(x)^{T}$ is symmetric positive definite. The first $d$ columns of $W$ consist of a basis of the tangent space $\mathcal{T}_{x}\mathcal{U}$ of $\mathcal{U}$ at $x$, and the remaining $m$ columns span the normal space $\mathcal{N}_{x}\mathcal{U}$. We then need to equip $\mathcal{Z}$ with the induced Riemannian metric
$d_{\text{{mod}}}(z,\tilde{z})=\inf_{\zeta(0)=z,\,\zeta(1)=\tilde{z}}\int_{0}^{1}\left(\dot{\zeta}(t)^{T}W^{-1}(\zeta(t))\dot{\zeta}(t)\right)^{1/2}\,\mathrm{d}t.$
The difficulty then usually consists in finding a more convenient form of $d_{\text{{mod}}}(z,\tilde{z})$, which allows to integrate the Fokker-Planck equation against $d_{\text{{mod}}}^{2}(z,\tilde{z})$ in order to show contraction. We leave this to future work.
Example 3.16.
We can check that in the one-dimensional quadratic case that was considered as an illustrative counterexample in Theorems 3.5 and 3.6, the modified equation now has the desired invariant distribution $\pi^{\ast}$. Furthermore, every solution $\pi_{t}$ of the Fokker-Planck equation converges to $\pi^{\ast}$ as $t\to\infty$. This can be seen as follows: We carry out the same computations as in the proof of Theorem 3.5 to obtain the Fourier transform $p_{t}$ of the solution $\pi_{t}$, given by
$p_{t}(\zeta)=p_{0}(\exp(-tA_{\text{{mod}}}^{T})\zeta)\exp\left(-\frac{1}{2}\int_{0}^{t}(B_{\text{{mod}}}^{T}\exp(-sA_{\text{{mod}}}^{T})\zeta)^{2}\,\mathrm{d}s\right),$
$A_{\text{{mod}}}=\begin{pmatrix}c_{g}^{-1}&k\\ -\lambda k+kc_{f}^{-1}c_{g}^{-1}+k^{3}c_{f}^{-2}&\lambda c_{f}\end{pmatrix},\quad B_{\text{{mod}}}=\begin{pmatrix}1\\ kc_{f}^{-1}\end{pmatrix},$
This time, the quadratic form under the integral is not positive definite in $\zeta$ since $B_{\text{{mod}}}$ is an eigenvector of $A_{\text{{mod}}}$ with eigenvalue $v_{h}\coloneqq c_{g}^{-1}+k^{2}c_{f}^{-1}$. Instead, it holds
$\int_{0}^{t}(B_{\text{{mod}}}^{T}\exp(-sA_{\text{{mod}}}^{T})\zeta)^{2}\,\mathrm{d}s=v_{h}^{-1}\left(1-e^{-tv_{h}}\right)(B_{\text{{mod}}}^{T}\zeta)^{2}.$
Taking the inverse Fourier transform shows that for every (possibly degenerate) initialization $\pi_{0}$ the solution $\pi_{t}$ converges to the invariant distribution, which is degenerated Gaussian with mean zero and covariance matrix $v_{h}^{-1}B_{\text{{mod}}}B_{\text{{mod}}}^{T}$. This is precisely the target distribution $\pi^{\ast}$.
4 Convergence of the Discrete Time Algorithm
We want to apply similar techniques as in Section 3 to the time discretized setting of the algorithm (1.4). Indeed, couplings like the ones we employed in the continuous time setting above have been used before in the stability and convergence analysis of Langevin algorithms. We begin this section by recalling such a computation in the simple case of ULA as an instructive example, and then expand to the primal-dual algorithm.
For a convex functional $h:\mathcal{X}\to\mathbb{R}\cup\{\infty\}$, points $x,\tilde{x}\in\mathcal{X}$ and subgradients $p\in\partial h(x),\tilde{p}\in\partial h(\tilde{x})$, we will use the symmetric Bregman distance $D_{h,\text{sym}}^{p,\tilde{p}}(x,\tilde{x})=\langle x-\tilde{x},p-\tilde{p}\rangle$. If $h$ is differentiable with unique subgradients, we will drop the reference to $p,\tilde{p}$ and just write $D_{h,\text{sym}}(x,\tilde{x})$. Recall that, if $h$ is convex, $D_{h,\text{sym}}$ is always non-negative.
Suppose now that the potential $h:\mathbb{R}^{d}\to\mathbb{R}$ is differentiable, that $\nabla h$ is $L_{\nabla h}$-Lipschitz-continuous and $\omega_{h}$-strongly convex. We consider two coupled Markov chains $(X_{n})_{n},(\tilde{X}_{n})_{n}$ defined by initial values $X_{0}\sim\mu_{0},\tilde{X}_{0}\sim\tilde{\mu}_{0}$ and
$\left\{\begin{array}[]{l}\xi^{n+1}\sim\mathrm{N}(0,I)\vspace{0.4em}\\ X^{n+1}=X^{n}-\tau\nabla h(X_{n})+\sqrt{2\tau}\xi^{n+1}\vspace{0.4em}\\ \tilde{X}^{n+1}=\tilde{X}^{n}-\tau\nabla h(\tilde{X}^{n})+\sqrt{2\tau}\xi^{n+1}\end{array}\right.$
Then we can compute that (see [Durmus2019], Prop. 3 where this calculation has been carried out)
$\lVert X^{n+1}-\tilde{X}^{n+1}\rVert^{2}$ $=\lVert X^{n}-\tau\nabla h(X^{n})-\tilde{X}^{n}+\tau\nabla h(\tilde{X}^{n})\rVert^{2}$
$=\lVert X^{n}-\tilde{X}^{n}\rVert^{2}-2\tau D_{h,\textrm{sym}}(X^{n},\tilde{X}^{n})+\tau^{2}\lVert\nabla h(X^{n})-\nabla h(\tilde{X}^{n})\rVert^{2}.$
If $h$ is $\omega_{h}$-strongly convex, we can use
$D_{h,\textrm{sym}}(X^{n},\tilde{X}^{n})\geq\frac{\omega_{h}L_{\nabla h}}{\omega_{h}+L_{\nabla h}}\lVert X^{n}-\tilde{X}^{n}\rVert^{2}+\frac{1}{\omega_{h}+L_{\nabla h}}\lVert\nabla h(X^{n})-\nabla h(\tilde{X}^{n})\rVert^{2},$
and we obtain
$\lVert X^{n+1}-\tilde{X}^{n+1}\rVert^{2}\leq\left(1-\tau\frac{2\omega_{h}L_{\nabla h}}{\omega_{h}+L_{\nabla h}}\right)\lVert X^{n}-\tilde{X}^{n}\rVert^{2}+\tau\left(\tau-\frac{2}{\omega_{h}+L_{\nabla h}}\right)\lVert\nabla h(X^{n})-\nabla h(\tilde{X}^{n})\rVert^{2}.$
If the step size is chosen to be $\tau=2/(\omega_{h}+L_{\nabla h})$, we obtain the convergence rate
$\lVert X^{n}-\tilde{X}^{n}\rVert^{2}\leq\left(\frac{L_{\nabla h}-\omega_{h}}{L_{\nabla h}+\omega_{h}}\right)^{2n}\lVert X^{0}-\tilde{X}^{0}\rVert^{2}.$
Denoting the distribution of $X^{n}$ by $\mu_{n}$ and that of $\tilde{X}^{n}$ by $\tilde{\mu}_{n}$, we obtain by taking the infimum over all couplings of $\mu_{0}$ and $\tilde{\mu}_{0}$ on the right side that
$\mathcal{W}_{2}(\mu_{n},\tilde{\mu}_{n})\leq\left(\frac{L_{\nabla h}-\omega_{h}}{L_{\nabla h}+\omega_{h}}\right)^{n}\mathcal{W}_{2}(\mu_{0},\tilde{\mu}_{0}).$
This shows that ULA is stable and will converge to its unique stationary distribution. Similar computations can be carried out for other time-discretized schemes based on Langevin diffusion, e.g. time-implicit leading to proximal sampling schemes. Since the stationary distribution of the time-discrete scheme is not the target distribution $\mu^{\ast}$, the missing step to relate the samples drawn from the Markov chain and the target is to couple a time-discrete chain and a time-continuous process and then finding an error bound. Typically, this is more challenging and the introduced bias (in Wasserstein distance) between the two settings scales with a power of the step size $\tau$.
4.1 Stability of Primal-Dual Langevin Sampling
We now turn to the primal-dual setting of (1.4). Consider the following coupling in discrete time with initial values $(X^{0},Y^{0})\sim\pi_{0}$ and $(\tilde{X}^{0},\tilde{Y}^{0})\sim\tilde{\pi}_{0}$ and identical stochastic input to both iterations:
$\left\{\begin{array}[]{l}\xi^{n+1}\sim\mathrm{N}(0,I)\vspace{0.8em}\\ Y^{n+1}=\operatorname{prox}_{\sigma f^{\ast}}(Y^{n}+\sigma KX_{\theta}^{n})\\ X^{n+1}=\operatorname{prox}_{\tau g}(X^{n}-\tau K^{T}Y^{n+1})+\sqrt{2\tau}\xi^{n+1}\\ X_{\theta}^{n+1}=X^{n+1}+\theta(X^{n+1}-X^{n})\vspace{0.8em}\\ \tilde{Y}^{n+1}=\operatorname{prox}_{\sigma f^{\ast}}(\tilde{Y}^{n}+\sigma K\tilde{X}_{\theta}^{n})\\ \tilde{X}^{n+1}=\operatorname{prox}_{\tau g}(\tilde{X}^{n}-\tau K^{T}\tilde{Y}^{n+1})+\sqrt{2\tau}\xi^{n+1}\\ \tilde{X}_{\theta}^{n+1}=\tilde{X}^{n+1}+\theta(\tilde{X}^{n+1}-\tilde{X}^{n})\end{array}\right.$ (4.1)
[…]
We start by giving a stability result for the case where neither $g$ nor $f^{\ast}$ are strongly convex.
Theorem 4.1.
Let $(X^{n},Y^{n})$ and $(\tilde{X}^{n},\tilde{Y}^{n})$ be given by the coupling (4.1) with relaxation parameter $\theta=1$ and $\tau\sigma L^{2}\leq 1$. Then for the primal variable it holds
$\mathcal{W}_{2}(\mu_{N},\tilde{\mu}_{N})\leq\mathcal{T}_{1,\lambda^{-1}}(\pi_{0},\tilde{\pi}_{0}),$
where $\mathcal{T}_{1,\lambda^{-1}}$ is the transport distance over the joint primal-dual variable defined in (3.1). If $\tau\sigma L^{2}<1$, then the joint variable obeys
$\mathcal{T}_{1,\lambda^{-1}}(\pi_{N},\tilde{\pi}_{N})\leq C\mathcal{T}_{1,\lambda^{-1}}(\pi_{0},\tilde{\pi}_{0}),$
with the constant $C=(1-\tau\sigma L^{2})^{-1/2}$.
In the case of strongly convex $g$ and $f^{\ast}$, we can prove the following contraction result.
Theorem 4.2.
Assume that $g$ and $f^{\ast}$ are $\omega_{g}$- and $\omega_{f^{\ast}}$-strongly convex respectively. Let $(X^{n},Y^{n})$ and $(\tilde{X}^{n},\tilde{Y}^{n})$ be given by the coupling (4.1) with extrapolation parameter $\theta$ satisfying
$\max\{(1+2\omega_{g}\tau)^{-1},(1+2\omega_{f^{\ast}}\sigma)^{-1}\}\leq\theta<1,$
and further $\theta\sigma\tau L^{2}\leq 1$. Then with the constant $\tilde{\lambda}\coloneqq\frac{1+2\omega_{g}\tau}{1+2\omega_{f^{\ast}}\sigma}\,\lambda$, it holds
$\mathcal{W}_{2}(\mu_{N},\tilde{\mu}_{N})\leq\theta^{N/2}\mathcal{T}_{1,\tilde{\lambda}^{-1}}(\pi_{0},\tilde{\pi}_{0}).$
and if $\theta\sigma\tau L^{2}<1$, then with $\tilde{C}\coloneqq(1-\theta\tau\sigma L^{2})^{-1/2}$ we have for the joint variable
$\mathcal{T}_{1,\tilde{\lambda}^{-1}}(\pi_{N},\tilde{\pi}_{N})\leq\tilde{C}\theta^{N/2}\mathcal{T}_{1,\tilde{\lambda}^{-1}}(\pi_{0},\tilde{\pi}_{0}).$
It follows directly that in the latter case a stationary distribution $\pi_{\lambda,\tau}$ exists and the distribution $\pi_{N}$ of samples converges to $\pi_{\lambda,\tau}$ as $N\to\infty$.
In order to prove Theorems 4.1 and 4.2, we need the following auxiliary inequality, which has been used several times in similar forms for primal-dual optimization algorithms, see e.g. [Chambolle2011]. It still holds in the present sampling setting, since considering the difference of two coupled processes makes the stochastic terms $\xi^{n}$ cancel out:
Lemma 4.3.
Let $(X^{n},Y^{n})$ and $(\tilde{X}^{n},\tilde{Y}^{n})$, $n\geq 0$ be given by the coupling (4.1) with extrapolation parameter $\theta$ chosen such that $\theta\sigma\tau L^{2}\leq 1$. By definition of the proximal mapping, we have for $n\geq 0$:
$\begin{array}[]{l}p^{n+1}\coloneqq(X^{n}-X^{n+1}+\sqrt{2\tau}\xi^{n+1})/\tau-K^{T}Y^{n+1},\in\partial g(X^{n+1}-\sqrt{2\tau}\xi^{n+1})\vspace{0.5em}\\ q^{n+1}\coloneqq(Y^{n}-Y^{n+1})/\sigma+KX_{\theta}^{n}\in\partial f^{\ast}(Y^{n+1}).\end{array}$ (4.2)
Analogously, we define $\tilde{p}^{n+1},\tilde{q}^{n+1}$ for the second chain $(\tilde{X}^{n},\tilde{Y}^{n})$. Then it holds
$2\left(D_{g,\text{sym}}^{p^{n+1},\tilde{p}^{n+1}}(X^{n+1}-\sqrt{2\tau}\xi^{n+1},\tilde{X}^{n+1}-\sqrt{2\tau}\xi^{n+1})+D_{f^{\ast},\text{sym}}^{q^{n+1},\tilde{q}^{n+1}}(Y^{n+1},\tilde{Y}^{n+1})\right)$
$\quad+\lVert X^{n+1}-\tilde{X}^{n+1}\rVert_{\tau^{-1}}^{2}+\lVert X^{n+1}-\tilde{X}^{n+1}-X^{n}+\tilde{X}^{n}\rVert_{\tau^{-1}}^{2}+\lVert Y^{n+1}-\tilde{Y}^{n+1}\rVert_{\sigma^{-1}}^{2}$
$\quad+2\langle K(X^{n+1}-\tilde{X}^{n+1}-X^{n}+\tilde{X}^{n}),Y^{n+1}-\tilde{Y}^{n+1}\rangle$
$\leq\lVert X^{n}-\tilde{X}^{n}\rVert_{\tau^{-1}}^{2}+\theta\lVert X^{n}-\tilde{X}^{n}-X^{n-1}+\tilde{X}^{n-1}\rVert_{\tau^{-1}}^{2}+\lVert Y^{n}-\tilde{Y}^{n}\rVert_{\sigma^{-1}}^{2}$
$\quad+2\theta\langle K(X^{n}-\tilde{X}^{n}-X^{n-1}+\tilde{X}^{n-1}),Y^{n}-\tilde{Y}^{n}\rangle.$
Proof.
In order to shorten the expressions, throughout the proof we abbreviate $U^{n}:=X^{n}-\tilde{X}^{n}$, $U_{\theta}^{n}=X_{\theta}^{n}-\tilde{X}_{\theta}^{n}$ and $V^{n}:=Y^{n}-\tilde{Y}^{n}$.
By plugging in the representation of the subgradient $p^{n+1}$ of $g$ at $X^{n+1}-\sqrt{2\tau}\xi^{n+1}$ (and likewise $\tilde{p}^{n+1}$) given in the statement of the lemma, we can rewrite a symmetric Bregman divergence as
$2D_{g,\text{sym}}^{p^{n+1},\tilde{p}^{n+1}}(X^{n+1}-\sqrt{2\tau}\xi^{n+1},\tilde{X}^{n+1}-\sqrt{2\tau}\xi^{n+1})$
$\quad=2\langle U^{n+1},p^{n+1}-\tilde{p}^{n+1}\rangle$
$\quad=2\langle U^{n+1},(X^{n}-X^{n+1})/\tau-K^{T}Y^{n+1}-(\tilde{X}^{n}-\tilde{X}^{n+1})/\tau+K^{T}\tilde{Y}^{n+1}\rangle$
$\quad=\frac{2}{\tau}\langle U^{n+1},U^{n}-U^{n+1}\rangle-2\langle U^{n+1},K^{T}V^{n+1}\rangle$
$\quad=\lVert U^{n}\rVert_{\tau^{-1}}^{2}-\lVert U^{n+1}\rVert_{\tau^{-1}}^{2}-\lVert U^{n+1}-U^{n}\rVert_{\tau^{-1}}^{2}-2\langle KU^{n+1},V^{n+1}\rangle,$
where we used the identity $-\langle A-B,A\rangle=\frac{1}{2}\lVert B\rVert^{2}-\frac{1}{2}\lVert A\rVert^{2}-\frac{1}{2}\lVert A-B\rVert^{2}$ in the last equality. Repeating the same calculation for the subgradients $q^{n+1}$ of $f^{\ast}$ at $Y^{n+1}$, we further get
$2D_{f^{\ast},\text{sym}}^{q^{n+1},\tilde{q}^{n+1}}(Y^{n+1},\tilde{Y}^{n+1})$ $=2\langle V^{n+1},(Y^{n}-Y^{n+1})/\sigma+KX_{\theta}^{n}-(\tilde{Y}^{n}-\tilde{Y}^{n+1})/\sigma-K\tilde{X}_{\theta}^{n}\rangle$
$=\frac{2}{\sigma}\langle V^{n+1},V^{n}-V^{n+1}\rangle+2\langle V^{n+1},KU_{\theta}^{n}\rangle$
$=\lVert V^{n}\rVert_{\sigma^{-1}}^{2}-\lVert V^{n+1}\rVert_{\sigma^{-1}}^{2}-\lVert V^{n+1}-V^{n}\rVert_{\sigma^{-1}}^{2}+2\langle KU_{\theta}^{n},V^{n+1}\rangle.$
Adding up the two equalities above gives
$2\langle U^{n+1},p^{n+1}-\tilde{p}^{n+1}\rangle+2\langle V^{n+1},q^{n+1}-\tilde{q}^{n+1}\rangle$ $=\lVert U^{n}\rVert_{\tau^{-1}}^{2}-\lVert U^{n+1}\rVert_{\tau^{-1}}^{2}-\lVert U^{n+1}-U^{n}\rVert_{\tau^{-1}}^{2}$
$\quad+\lVert V^{n}\rVert_{\sigma^{-1}}^{2}-\lVert V^{n+1}\rVert_{\sigma^{-1}}^{2}-\lVert V^{n+1}-V^{n}\rVert_{\sigma^{-1}}^{2}$
$\quad-2\langle KU^{n+1},V^{n+1}\rangle+2\langle KU_{\theta}^{n},V^{n+1}\rangle.$ (4.3)
By definition it holds $U_{\theta}^{n}=U^{n}+\theta(U^{n}-U^{n-1})$, so the two inner products in the last line can be rewritten in the following way
$-2\langle KU^{n+1},V^{n+1}\rangle+2\langle KU_{\theta}^{n},V^{n+1}\rangle$
$=-2\langle K(U^{n+1}-U^{n}),V^{n+1}\rangle+2\theta\langle K(U^{n}-U^{n-1}),V^{n+1}\rangle$
$=-2\langle K(U^{n+1}-U^{n}),V^{n+1}\rangle+2\theta\langle K(U^{n}-U^{n-1}),V^{n}\rangle+2\theta\langle K(U^{n}-U^{n-1}),V^{n+1}-V^{n}\rangle$
$\leq-2\langle K(U^{n+1}-U^{n}),V^{n+1}\rangle+2\theta\langle K(U^{n}-U^{n-1}),V^{n}\rangle$
$\quad+\theta\lVert U^{n}-U^{n-1}\rVert_{\tau^{-1}}^{2}+\tau\sigma\theta L^{2}\lVert V^{n+1}-V^{n}\rVert_{\sigma^{-1}}^{2}.$ (4.4)
Inserting this into (4.1) and using $(1-\sigma\tau\theta L^{2})\lVert V^{n+1}-V^{n}\rVert_{\sigma^{-1}}^{2}\geq 0$ gives the statement of the Lemma. ∎
Proof of Theorem 4.1.
In the convex case, we start by noting that symmetric Bregman distances for convex functions are always non-negative, i.e.
$D_{g,\text{sym}}^{p^{n+1},\tilde{p}^{n+1}}(X^{n+1}-\sqrt{2\tau}\xi^{n+1},\tilde{X}^{n+1}-\sqrt{2\tau}\xi^{n+1})\geq 0,$
$D_{f^{\ast},\text{sym}}^{q^{n+1},\tilde{q}^{n+1}}(Y^{n+1},\tilde{Y}^{n+1})\geq 0.$
Setting $\theta=1$, the inequality in Lemma 4.3 hence gives
$\lVert X^{n+1}-\tilde{X}^{n+1}\rVert_{\tau^{-1}}^{2}+\lVert X^{n+1}-\tilde{X}^{n+1}-X^{n}+\tilde{X}^{n}\rVert_{\tau^{-1}}^{2}+\lVert Y^{n+1}-\tilde{Y}^{n+1}\rVert_{\sigma^{-1}}^{2}$
$\quad+2\langle K(X^{n+1}-\tilde{X}^{n+1}-X^{n}+\tilde{X}^{n}),Y^{n+1}-\tilde{Y}^{n+1}\rangle+(1-\sigma\tau L^{2})\lVert Y^{n+1}-\tilde{Y}^{n+1}-Y^{n}+\tilde{Y}^{n}\rVert_{\sigma^{-1}}^{2}$
$\leq\lVert X^{n}-\tilde{X}^{n}\rVert_{\tau^{-1}}^{2}+\lVert X^{n}-\tilde{X}^{n}-X^{n-1}+\tilde{X}^{n-1}\rVert_{\tau^{-1}}^{2}+\lVert Y^{n}-\tilde{Y}^{n}\rVert_{\sigma^{-1}}^{2}$
$\quad+2\langle K(X^{n}-\tilde{X}^{n}-X^{n-1}+\tilde{X}^{n-1}),Y^{n}-\tilde{Y}^{n}\rangle.$
Iterating this inequality over $n=0,\dots,N-1$ steps and using $X^{-1}=X^{0}$, $\tilde{X}^{-1}=\tilde{X}^{0}$ gives
$\lVert X^{N}-\tilde{X}^{N}\rVert_{\tau^{-1}}^{2}+\lVert Y^{N}-\tilde{Y}^{N}\rVert_{\sigma^{-1}}^{2}+\lVert X^{N}-\tilde{X}^{N}-X^{N-1}+\tilde{X}^{N-1}\rVert_{\tau^{-1}}^{2}$
$\quad+2\langle K(X^{N}-\tilde{X}^{N}-X^{N-1}+\tilde{X}^{N-1}),Y^{N}-\tilde{Y}^{N}\rangle$
$\leq\lVert X^{0}-\tilde{X}^{0}\rVert_{\tau^{-1}}^{2}+\lVert Y^{0}-\tilde{Y}^{0}\rVert_{\sigma^{-1}}^{2}$
Using the bound $2\langle a,b\rangle\geq-c^{-1}\lVert a\rVert^{2}-c\lVert b\rVert^{2}$ with $c=\tau L^{2}$ on the remaining inner product implies
$\lVert X^{N}-\tilde{X}^{N}\rVert_{\tau^{-1}}^{2}+(1-\tau\sigma L^{2})\lVert Y^{N}-\tilde{Y}^{N}\rVert_{\sigma^{-1}}^{2}\leq\lVert X^{0}-\tilde{X}^{0}\rVert_{\tau^{-1}}^{2}+\lVert Y^{0}-\tilde{Y}^{0}\rVert_{\sigma^{-1}}^{2}.$
Taking expectations on both sides, by definition of the Wasserstein distance, the left side is lower bounded by $\tau^{-1}\mathcal{W}_{2}^{2}(\mu_{N},\tilde{\mu}_{N})$. Since the resulting inequality holds for all couplings of initializations $(X^{0},Y^{0})$ and $(\tilde{X}^{0},\tilde{Y}^{0})$, taking the infimum over all couplings on the right side gives $\tau^{-1}\mathcal{T}^{2}_{1,\lambda^{-1}}(\pi_{0},\tilde{\pi}_{0})$, proving the first inequality. The second inequality in the statement of Theorem 4.1 results from lower bounding the left side not by the Wasserstein distance in the primal variable, but instead by $C^{-2}\tau^{-1}\mathcal{T}^{2}_{1,\lambda^{-1}}(\pi_{N},\tilde{\pi}_{N})$. ∎
Proof of Theorem 4.2.
We now assume strong convexity of $g,f^{\ast}$. In that case, the symmetric Bregman distances obey the following lower bounds
$D_{g,\text{sym}}^{p^{n+1},\tilde{p}^{n+1}}(X^{n+1}-\sqrt{2\tau}\xi^{n+1},\tilde{X}^{n+1}-\sqrt{2\tau}\xi^{n+1})\geq\omega_{g}\lVert X^{n+1}-\tilde{X}^{n+1}\rVert^{2},$
$D_{f^{\ast},\text{sym}}^{q^{n+1},\tilde{q}^{n+1}}(Y^{n+1},\tilde{Y}^{n+1})\geq\omega_{f^{\ast}}\lVert Y^{n+1}-\tilde{Y}^{n+1}\rVert^{2}.$
The inequality in Lemma 4.3 now implies
$\lVert X^{n+1}-\tilde{X}^{n+1}\rVert_{\tau^{-1}+2\omega_{g}}^{2}+\lVert X^{n+1}-\tilde{X}^{n+1}-X^{n}+\tilde{X}^{n}\rVert_{\tau^{-1}}^{2}+\lVert Y^{n+1}-\tilde{Y}^{n+1}\rVert_{\sigma^{-1}+2\omega_{f^{\ast}}}^{2}$
$\quad+2\langle K(X^{n+1}-\tilde{X}^{n+1}-X^{n}+\tilde{X}^{n}),Y^{n+1}-\tilde{Y}^{n+1}\rangle$
$\leq\lVert X^{n}-\tilde{X}^{n}\rVert_{\tau^{-1}}^{2}+\theta\lVert X^{n}-\tilde{X}^{n}-X^{n-1}+\tilde{X}^{n-1}\rVert_{\tau^{-1}}^{2}+\lVert Y^{n}-\tilde{Y}^{n}\rVert_{\sigma^{-1}}^{2}$
$\quad+2\theta\langle K(X^{n}-\tilde{X}^{n}-X^{n-1}+\tilde{X}^{n-1}),Y^{n}-\tilde{Y}^{n}\rangle.$
Define now
$\Delta^{n}$ $\coloneqq\lVert X^{n}-\tilde{X}^{n}\rVert_{\tau^{-1}+2\omega_{g}}^{2}+\lVert X^{n}-\tilde{X}^{n}-X^{n-1}+\tilde{X}^{n-1}\rVert_{\tau^{-1}}^{2}+\lVert Y^{n}-\tilde{Y}^{n}\rVert_{\sigma^{-1}+2\omega_{f^{\ast}}}^{2}$
$\quad+2\langle K(X^{n}-\tilde{X}^{n}-X^{n-1}+\tilde{X}^{n-1}),Y^{n}-\tilde{Y}^{n}\rangle.$
Then under the assumptions on $\theta$ given in Theorem 4.2, it holds
$\Delta^{n+1}$ $\leq\lVert X^{n}-\tilde{X}^{n}\rVert_{\tau^{-1}}^{2}+\theta\lVert X^{n}-\tilde{X}^{n}-X^{n-1}+\tilde{X}^{n-1}\rVert_{\tau^{-1}}^{2}+\lVert Y^{n}-\tilde{Y}^{n}\rVert_{\sigma^{-1}}^{2}$
$\quad+2\theta\langle K(X^{n}-\tilde{X}^{n}-X^{n-1}+\tilde{X}^{n-1}),Y^{n}-\tilde{Y}^{n}\rangle$
$\leq\theta\Delta^{n}$
Iterating this inequality over $n=0,\dots,N-1$ we obtain
$\Delta^{N}$ $\leq\theta^{N}\Delta^{0}.$
We use again Cauchy-Schwarz and Young’s inequality on the inner product in $\Delta^{N}$ as well as the initial condition $X^{-1}=X^{0}$, $\tilde{X}^{-1}=X^{0}$ and obtain
$\lVert X^{N}-\tilde{X}^{N}\rVert_{\tau^{-1}+2\omega_{g}}^{2}+(1-\theta\tau\sigma L^{2})\lVert Y^{N}-\tilde{Y}^{N}\rVert_{\sigma^{-1}+2\omega_{f^{\ast}}}^{2}$
$\quad\leq\theta^{N}\left(\lVert X^{0}-\tilde{X}^{0}\rVert_{\tau^{-1}+2\omega_{g}}^{2}+\lVert Y^{0}-\tilde{Y}^{0}\rVert_{\sigma^{-1}+2\omega_{f^{\ast}}}^{2}\right)$
From here, we proceed as in the proof of Theorem 4.1: Taking expectations on both sides, the left side can be bounded below by $(\tau^{-1}+2\omega_{g})W_{2}^{2}(\mu_{N},\tilde{\mu}_{N})$ or alternatively in the joint variable by $\tilde{C}^{-2}(\tau^{-1}+2\omega_{g})\mathcal{T}_{1,\tilde{\lambda}^{-1}}^{2}(\pi_{N},\tilde{\pi}_{N})$. The transport distance on the right is obtained by taking the infimum over all initialization couplings, completing the proof. ∎
Having proven the stability theorems for the iteration (1.4), we briefly comment on the version of the algorithm that adds the stochastic term $\xi$ inside the argument of the proximal operator.
Corollary 4.4.
Instead of (1.4), consider the iteration procedure
$\left\{\begin{array}[]{l}\xi^{n+1}\sim\mathrm{N}(0,I)\\ Y^{n+1}=\operatorname{prox}_{\sigma f^{\ast}}(Y^{n}+\sigma KX_{\theta}^{n})\\ X^{n+1}=\operatorname{prox}_{\tau g}(X^{n}-\tau K^{T}Y^{n+1}+\sqrt{2\tau}\xi^{n+1})\\ X_{\theta}^{n+1}=X^{n+1}+\theta(X^{n+1}-X^{n})\end{array}\right.$ (4.5)
with the stochastic term $\sqrt{2\tau}\xi$ added inside the proximal mapping. For the iterates of this algorithm, under the same assumptions, the results of Theorems 4.1 and 4.2 also hold. The stationary distribution of this modified algorithm, however, may differ from the stationary state $\pi_{\lambda,\tau}$ of (1.4).
Proof.
This can be seen by checking that the central inequality in Lemma 4.3 still holds. Under the changed iteration, we have
$\begin{array}[]{l}p^{n+1}\coloneqq(X^{n}-X^{n+1}+\sqrt{2\tau}\xi^{n+1})/\tau-K^{T}Y^{n+1},\in\partial g(X^{n+1}).\end{array}$
Note the difference to (4.2), where $p^{n+1}$ was a subgradient at the point $X^{n+1}-\sqrt{2\tau}\xi^{n+1}$ instead. The only change we need to make in the proof of Lemma 4.3 is that instead of the symmetric Bregman distance at the point $(X^{n+1}-\sqrt{2\tau}\xi^{n+1},\tilde{X}^{n+1}-\sqrt{2\tau}\xi^{n+1})$, we now start with
$D_{g,\text{sym}}^{p^{n+1},\tilde{p}^{n+1}}(X^{n+1},\tilde{X}^{n+1})=\langle U^{n+1},p^{n+1}-\tilde{p}^{n+1}\rangle.$
All the following computations remain unchanged. ∎
4.2 Convergence to the Continuous Time Stationary Solution
Having shown that the algorithm is stable with a limiting stationary distribution $\pi_{\lambda,\tau}$ in the strongly convex case, we need to find bounds on the distance between $\pi_{\lambda,\tau}$ and the continuous time stationary solution $\pi_{\lambda}$, in order to relate the drawn samples to the target. We approach this by considering again a coupling, but this time between the algorithm (1.4) and the stationary state of its continuous time limiting SDE (2.6).
Assume that the time steps $\tau,\sigma$ are constant with $\lambda=\sigma/\tau$. Let $(X^{0},Y^{0})\sim\pi^{0}$ for some initial distribution $\pi^{0}$, and set $X_{\theta}^{0}=X^{0}$. Now consider a process $\tilde{Z}_{t}=(\tilde{X}_{t},\tilde{Y}_{t})$, $t\in\mathbb{R}_{\geq 0}$ which is initialized at $\tilde{Z}_{0}\sim\pi_{\lambda}$. Let $W_{t}$ be $\mathbb{R}^{d}$-valued Brownian motion and denote $\xi^{n+1}:=\tau^{-1/2}\int_{n\tau}^{(n+1)\tau}\mathrm{d}W_{t}$ for all $n\in\mathbb{N}_{0}$. We now consider $Z^{n}=(X^{n},Y^{n})$, $n\in\mathbb{N}_{0}$ and $\tilde{Z}_{t}=(\tilde{X}_{t},\tilde{Y}_{t})$, $t\in\mathbb{R}_{\geq 0}$ defined by the coupling
$\left\{\begin{array}[]{l}Y^{n+1}=\operatorname{prox}_{\sigma f^{\ast}}(Y^{n}+\sigma KX_{\theta}^{n})\\ X^{n+1}=\operatorname{prox}_{\tau g}(X^{n}-\tau K^{T}Y^{n+1})+\sqrt{2\tau}\xi^{n+1}\\ X_{\theta}^{n+1}=X^{n+1}+\theta(X^{n+1}-X^{n})\vspace{0.8em}\\ \,\mathrm{d}\tilde{X}_{t}=-(\nabla g(\tilde{X}_{t})+K^{T}\tilde{Y}_{t})\,\mathrm{d}t+\sqrt{2}\,\mathrm{d}W_{t}\\ \,\mathrm{d}\tilde{Y}_{t}=-\lambda(\nabla f^{\ast}(\tilde{Y}_{t})-K\tilde{X}_{t})\,\mathrm{d}t.\end{array}\right.$ (4.6)
It immediately follows that the marginal distribution of $\tilde{Z}_{t}$ is stationary with $\tilde{Z}_{t}\sim\pi_{\lambda}$, $t\in\mathbb{R}_{\geq 0}$. For notational convenience, we also define an overrelaxation variable in continuous time by $\tilde{X}_{t,\theta}:=\tilde{X}_{t}+\theta(\tilde{X}_{t}-\tilde{X}_{t-\tau})$ and set $X_{t-\tau}=X_{0}$ when $t\in[0,\tau)$. For the coupling (4.6), we can prove an inequality similar to Lemma 4.3, but with additional discretization errors.
Lemma 4.5.
Let $(X^{n},Y^{n})$, $n\geq 1$, and $(X_{t},Y_{t})$, $t\geq 0$ be given by (4.6) with $\theta\sigma\tau L^{2}\leq 1$. The subgradients $p^{n}\in\partial g(X^{n}-\sqrt{2\tau}\xi^{n})$ and $q^{n}\in\partial f^{\ast}(Y^{n})$ at the discrete time iterates are given in (4.2). Fix now some $n\geq 0$, and let $t=n\tau$. Then we have the following inequality.
$2\left(D_{g,\text{sym}}^{p^{n+1},\tilde{p}^{\prime}}(X^{n+1}-\sqrt{2\tau}\xi^{n+1},\tilde{X}_{t+\tau}-\sqrt{2\tau}\xi^{n+1})+D_{f^{\ast},\text{sym}}^{q^{n+1},\tilde{q}_{t+\tau}}(Y^{n+1},\tilde{Y}_{t+\tau})\right)$
$\quad+\lVert X^{n+1}-\tilde{X}_{t+\tau}\rVert_{\tau^{-1}}^{2}+\lVert X^{n+1}-\tilde{X}_{t+\tau}-X^{n}+\tilde{X}_{t}\rVert_{\tau^{-1}}^{2}+\lVert Y^{n+1}-\tilde{Y}_{t+\tau}\rVert_{\sigma^{-1}}^{2}$
$\quad+2\langle K(X^{n+1}-\tilde{X}_{t+\tau}-X^{n}+\tilde{X}_{t}),Y^{n+1}-\tilde{Y}_{t+\tau}\rangle$
$\leq\lVert X^{n}-\tilde{X}_{t}\rVert_{\tau^{-1}}^{2}+\theta\lVert X^{n}-\tilde{X}_{t}-X^{n-1}+\tilde{X}_{t-\tau}\rVert_{\tau^{-1}}^{2}+\lVert Y^{n}-\tilde{Y}_{t}\rVert_{\sigma^{-1}}^{2}$
$\quad+2\theta\langle K(X^{n}-\tilde{X}_{t}-X^{n-1}+\tilde{X}_{t-\tau}),Y^{n}-\tilde{Y}_{t}\rangle$
$\quad+\frac{2}{\tau}\langle X^{n+1}-\tilde{X}_{t+\tau},\int_{t}^{t+\tau}(\tilde{p}_{s}-p^{\prime}+K^{T}(\tilde{Y}_{s}-\tilde{Y}_{t+\tau}))\,\mathrm{d}s\rangle$
$\quad+\frac{2}{\tau}\langle Y^{n+1}-\tilde{Y}_{t+\tau},\int_{t}^{t+\tau}(\tilde{q}_{s}-\tilde{q}_{t+\tau}-K(\tilde{X}_{s}-\tilde{X}_{t,\theta}))\,\mathrm{d}s\rangle,$
where $p^{\prime}\in\partial g(\tilde{X}_{t+\tau}-\sqrt{2\tau}\xi^{n+1})$ arbitrary.
Proof.
As before, we introduce the abbreviations $U^{n}:=X^{n}-\tilde{X}_{t}$, $U_{\theta}^{n}=X_{\theta}^{n}-\tilde{X}_{t,\theta}$ and $V^{n}:=Y^{n}-\tilde{Y}_{t}$.
Integrating the SDE for the continuous time primal variable $\tilde{X}$ over $[t,t+\tau]$, we obtain the following identity
$\tilde{X}_{t+\tau}=\tilde{X}_{t}-\int_{t}^{t+\tau}(\tilde{p}_{s}+K^{T}\tilde{Y}_{s})\,\mathrm{d}s+\sqrt{2\tau}\xi^{n+1}$
$\Leftrightarrow\frac{1}{\tau}\int_{t}^{t+\tau}\tilde{p}_{s}\,\mathrm{d}s=(\tilde{X}_{t}-\tilde{X}_{t+\tau}+\sqrt{2\tau}\xi^{n+1})/\tau-\frac{1}{\tau}\int_{t}^{t+\tau}K^{T}\tilde{Y}_{s}\,\mathrm{d}s.$ (4.7)
As in the proof of Lemma 4.3, we now rewrite the symmetric Bregman distance of $g$ at the point $(X^{n+1}-\sqrt{2\tau}\xi^{n+1},\tilde{X}_{t+\tau}-\sqrt{2\tau}\xi^{n+1})$. As before, we insert the representation of the subgradient $p^{n+1}$ of $g$ at $X^{n+1}-\sqrt{2\tau}\xi^{n+1}$. For the continuous time term, we add zero by adding and subtracting the two sides of (4.7) in order to separate the discretization error in the drift:
$2D_{g,\text{sym}}^{p^{n+1},p^{\prime}}(X^{n+1}-\sqrt{2\tau}\xi^{n+1},\tilde{X}_{t+\tau}-\sqrt{2\tau}\xi^{n+1})$
$=2\langle U^{n+1},p^{n+1}-p^{\prime}\rangle$
$=2\langle U^{n+1},(X^{n}-X^{n+1})/\tau-K^{T}Y^{n+1}-(\tilde{X}_{t}-\tilde{X}_{t+\tau})/\tau+\frac{1}{\tau}\int_{t}^{t+\tau}K^{T}\tilde{Y}_{s}\,\mathrm{d}s+\frac{1}{\tau}\int_{t}^{t+\tau}(\tilde{p}_{s}-p^{\prime})\,\mathrm{d}s\rangle$
$=\frac{2}{\tau}\langle U^{n+1},U^{n}-U^{n+1}\rangle-2\langle U^{n+1},K^{T}V^{n+1}\rangle+\frac{2}{\tau}\langle U^{n+1},\int_{t}^{t+\tau}(\tilde{p}_{s}-p^{\prime}+K^{T}(\tilde{Y}_{s}-\tilde{Y}_{t+\tau}))\,\mathrm{d}s\rangle$
$=\lVert U^{n}\rVert_{\tau^{-1}}^{2}-\lVert U^{n+1}\rVert_{\tau^{-1}}^{2}-\lVert U^{n+1}-U^{n}\rVert_{\tau^{-1}}^{2}-2\langle KU^{n+1},V^{n+1}\rangle$
$\quad+\frac{2}{\tau}\langle U^{n+1},\int_{t}^{t+\tau}(\tilde{p}_{s}-p^{\prime}+K^{T}(\tilde{Y}_{s}-\tilde{Y}_{t+\tau}))\,\mathrm{d}s\rangle.$
Similarly, for the dual variable in continuous time, we obtain the integrated SDE
$\tilde{Y}_{t+\tau}=\tilde{Y}_{t}-\lambda\int_{t}^{t+\tau}(\tilde{q}_{s}-K\tilde{X}_{s})\,\mathrm{d}s$
$\Leftrightarrow\quad\frac{1}{\tau}\int_{t}^{t+\tau}\tilde{q}_{s}\,\mathrm{d}s=(\tilde{Y}_{t}-\tilde{Y}_{t+\tau})/\sigma+\frac{1}{\tau}\int_{t}^{t+\tau}K\tilde{X}_{s}\,\mathrm{d}s.$ (4.8)
By plugging in $q^{n+1}$ from (4.2) and adding and subtracting the sides of (4.8), we can rewrite the symmetric Bregman distance of $f^{\ast}$ at $(Y^{n+1},\tilde{Y}_{t+\tau})$ as
$2\langle Y^{n+1}-\tilde{Y}_{t+\tau},q^{n+1}-\tilde{q}_{t+\tau}\rangle$
$=2\langle V^{n+1},(Y^{n}-Y^{n+1})/\sigma+KX^{n}_{\theta}-(\tilde{Y}_{t}-\tilde{Y}_{t+\tau})/\sigma-\frac{1}{\tau}\int_{t}^{t+\tau}K\tilde{X}_{s}\,\mathrm{d}s+\frac{1}{\tau}\int_{t}^{t+\tau}(\tilde{q}_{s}-\tilde{q}_{t+\tau})\,\mathrm{d}s\rangle$
$=\frac{2}{\sigma}\langle V^{n+1},V^{n}-V^{n+1}\rangle+2\langle V^{n+1},KU^{n}_{\theta}\rangle+\frac{2}{\tau}\langle V^{n+1},\int_{t}^{t+\tau}(\tilde{q}_{s}-\tilde{q}_{t+\tau}-K(\tilde{X}_{s}-\tilde{X}_{t,\theta}))\,\mathrm{d}s\rangle$
$=\lVert V^{n}\rVert_{\sigma^{-1}}^{2}-\lVert V^{n+1}\rVert_{\sigma^{-1}}^{2}-\lVert V^{n+1}-V^{n}\rVert_{\sigma^{-1}}^{2}+2\langle V^{n+1},KU^{n}_{\theta}\rangle$
$\quad+\frac{2}{\tau}\langle V^{n+1},\int_{t}^{t+\tau}(\tilde{q}_{s}-\tilde{q}_{t+\tau}-K(\tilde{X}_{s}-\tilde{X}_{t,\theta}))\,\mathrm{d}s\rangle$
Adding the primal and dual Bregman distances, we obtain
22 22 $\langle U^{n+1},p^{n+1}-p^{\prime}\rangle+2\langle V^{n+1},q^{n+1}-\tilde{q}^{n+1}\rangle$
$=\lVert U^{n}\rVert_{\tau^{-1}}^{2}-\lVert U^{n+1}\rVert_{\tau^{-1}}^{2}-\lVert U^{n+1}-U^{n}\rVert_{\tau^{-1}}^{2}-2\langle KU^{n+1},V^{n+1}\rangle$
$\quad+\lVert V^{n}\rVert_{\sigma^{-1}}^{2}-\lVert V^{n+1}\rVert_{\sigma^{-1}}^{2}-\lVert V^{n+1}-V^{n}\rVert_{\sigma^{-1}}^{2}+2\langle KU_{\theta}^{n},V^{n+1}\rangle$
$\quad+\frac{2}{\tau}\langle U^{n+1},\int_{t}^{t+\tau}(\tilde{p}_{s}-p^{\prime}+K^{T}(\tilde{Y}_{s}-\tilde{Y}_{t+\tau}))\,\mathrm{d}s\rangle$
$\quad+\frac{2}{\tau}\langle V^{n+1},\int_{t}^{t+\tau}(\tilde{q}_{s}-\tilde{q}_{t+\tau}-K(\tilde{X}_{s}-\tilde{X}_{t,\theta}))\,\mathrm{d}s\rangle.$
As in the proof of Lemma 4.3, the argument is complete by plugging in inequality (4.4). ∎
The main difference between Lemma 4.3 and Lemma 4.5 are the additional two terms due to the discretization error. The next auxiliary result provides an upper bound to these errors, controlled by the primal step size $\tau$.
Lemma 4.6.
Assume that $g\in C^{1}(\mathcal{X})$ and $\nabla g$ is Lipschitz continuous with constant $L_{\nabla g}>0$, and that $f$ is $\omega_{f}$-strongly convex with $\omega_{f}>0$. Assume further that there exists a constant $D_{4}$ such that
$\mathbb{E}_{Z\sim\pi_{\lambda}}\left[\lVert A_{\text{{PD}}}(Z)\rVert^{2}\right]\leq D_{4}.$
Let $Z_{0}\sim\pi_{\lambda}$ and $Z_{t}=(X_{t},Y_{t}),t>0$ satisfy the SDE (2.6), with $p_{t}\in\nabla g(X_{t})$ and $q_{t}\in\nabla f^{\ast}(Y_{t})$. For some $\tau>0$ and $n\in\mathbb{N}_{0}$, fix $t=n\tau$. Let further $p^{\prime}\in\partial g(X_{t+\tau}-\sqrt{2\tau}\xi)$ arbitrary, where $\xi=\tau^{-1/2}\int_{t}^{t+\tau}\mathrm{d}W_{s}$, and denote $X_{t,\theta}$ as before. Then we have the following bounds
$\mathbb{E}\left[\bigg{\lVert}\int_{t}^{t+\tau}(p_{s}-p^{\prime}-K^{T}(Y_{s}-Y_{t+\tau}))\,\mathrm{d}s\bigg{\rVert}^{2}\right]\leq 4L_{\nabla g}^{2}d\tau^{3}+\mathcal{O}(\tau^{4}),$
$\mathbb{E}\left[\bigg{\lVert}\int_{t}^{t+\tau}(q_{s}-q_{t+\tau}+K(X_{s}-X_{t,\theta}))\,\mathrm{d}s\bigg{\rVert}^{2}\right]\leq 6L^{2}(1+2\theta^{2})d\tau^{3}+\mathcal{O}(\tau^{4}).$
Proof.
We start with the first inequality. Applying first the Cauchy-Schwarz inequality and then the definitions of matrix norm and Lipschitz continuity, we can estimate
$\mathbb{E}\left[\bigg{\lVert}\int_{t}^{t+\tau}(p_{s}-p^{\prime}-K^{T}(Y_{s}-Y_{t+\tau}))\,\mathrm{d}s\bigg{\rVert}^{2}\right]$
$\leq 2\tau\int_{t}^{t+\tau}\mathbb{E}\left[\lVert p_{s}-p^{\prime}\rVert^{2}\right]\,\mathrm{d}s+2\tau\int_{t}^{t+\tau}\mathbb{E}\left[\lVert K^{T}(Y_{s}-Y_{t+\tau})\rVert^{2}\right]\,\mathrm{d}s$
$\leq 2\tau L_{\nabla g}^{2}\int_{t}^{t+\tau}\mathbb{E}\left[\lVert X_{s}-X_{t+\tau}+\sqrt{2\tau}\xi\rVert^{2}\right]\,\mathrm{d}s+2\tau L^{2}\int_{t}^{t+\tau}\mathbb{E}\left[\lVert Y_{s}-Y_{t+\tau}\rVert^{2}\right]\,\mathrm{d}s$ (4.9)
where $L_{\nabla g}$ is the Lipschitz constant of $\nabla g$ and $L=\lVert K\rVert$. We now plug in the SDE, integrated over $[s,t+\tau]$, for both $X_{t+\tau}$ and $Y_{t+\tau}$ and obtain
$X_{s}-X_{t+\tau}+\sqrt{2\tau}\xi$ $=\int_{s}^{t+\tau}(p_{r}+K^{T}Y_{r})\,\mathrm{d}r-\sqrt{2}\int_{t}^{s}\mathrm{d}W_{r}$
$Y_{s}-Y_{t+\tau}$ $=\int_{s}^{t+\tau}(q_{r}-KX_{r})\,\mathrm{d}r.$
The continuous time process is stationary in distribution at $(X_{r},Y_{r})\sim\pi_{\lambda}$ for all $r>0$. Hence, by assumption we have $\mathbb{E}\left[\lVert p_{r}+K^{T}Y_{r}\rVert^{2}\right]\leq M$ and $\mathbb{E}\left[\lVert q_{r}-KX_{r}\rVert^{2}\right]\leq M$. We can therefore estimate for all $t\leq s\leq t+\tau$:
$\mathbb{E}\left[\lVert X_{s}-X_{t+\tau}+\sqrt{2\tau}\xi\rVert^{2}\right]$ $\leq 2\mathbb{E}\left[\bigg{\lVert}\int_{s}^{t+\tau}(p_{r}+K^{T}Y_{r})\,\mathrm{d}r\bigg{\rVert}^{2}\right]+4\mathbb{E}\left[\bigg{\lVert}\int_{t}^{s}\mathrm{d}W_{r}\bigg{\rVert}^{2}\right]$
$\leq 2(t+\tau-s)\int_{s}^{t+\tau}\mathbb{E}\left[\lVert p_{r}+K^{T}Y_{r}\rVert^{2}\right]\,\mathrm{d}r+4d(s-t)$
$\leq 2D_{4}(t+\tau-s)^{2}+4d(s-t)$
$\mathbb{E}\left[\lVert Y_{s}-Y_{t+\tau}\rVert^{2}\right]$ $=\mathbb{E}\left[\bigg{\lVert}\int_{s}^{t+\tau}(q_{r}-KX_{r})\,\mathrm{d}r\bigg{\rVert}^{2}\right]$
$\leq D_{4}(t+\tau-s)^{2}$ (4.10)
Plugging these two bounds into (4.9) and integrating gives the desired bound
$\mathbb{E}\left[\bigg{\lVert}\int_{t}^{t+\tau}(p_{s}-p^{\prime}-K^{T}(Y_{s}-Y_{t+\tau}))\,\mathrm{d}s\bigg{\rVert}^{2}\right]\leq 4L_{\nabla g}^{2}d\tau^{3}+\frac{2D_{4}(2L_{\nabla g}^{2}+L^{2})}{3}\tau^{4}.$
The bound for the discretization error in dual space can be derived using similar estimates
$\mathbb{E}\left[\bigg{\lVert}\int_{t}^{t+\tau}(q_{s}-q_{t+\tau}+K(X_{s}-X_{t,\theta}))\,\mathrm{d}s\bigg{\rVert}^{2}\right]$
$\leq 2\tau\int_{t}^{t+\tau}\mathbb{E}\left[\lVert q_{s}-q_{t+\tau}\rVert^{2}\right]\,\mathrm{d}s+\int_{t}^{t+\tau}\mathbb{E}\left[\lVert K(X_{s}-X_{t,\theta})\rVert^{2}\right]\,\mathrm{d}s$
$\leq\frac{2\tau}{\omega_{f}^{2}}\int_{t}^{t+\tau}\mathbb{E}\left[\lVert Y_{s}-Y_{t+\tau}\rVert^{2}\right]\,\mathrm{d}s+2\tau L^{2}\int_{t}^{t+\tau}\mathbb{E}\left[\lVert X_{s}-X_{t,\theta}\rVert^{2}\right]\,\mathrm{d}s$ (4.11)
If $n\geq 1$, the term in the second integral can be bounded using
$X_{s}-X_{t,\theta}$ $=X_{s}-X_{t}-\theta(X_{t}-X_{t-\tau})$
$=-\int_{t}^{s}(p_{r}+K^{T}Y_{r})\,\mathrm{d}r+\sqrt{2}\int_{t}^{s}\mathrm{d}W_{r}+\theta\int_{t-\tau}^{t}(p_{r}+K^{T}Y_{r})\,\mathrm{d}r-\sqrt{2}\theta\int_{t-\tau}^{t}\mathrm{d}W_{r}.$
In the boundary case $n=0$ implying $t=0$, we have $X_{t-\tau}=X_{t}=X_{0}$ and the last two integrals can be dropped. In either case, we have the following bound
$\mathbb{E}\left[\lVert X_{s}-X_{t,\theta}\rVert^{2}\right]$ $\leq 3\mathbb{E}\left[\bigg{\lVert}\int_{t}^{s}(p_{r}+K^{T}Y_{r})\,\mathrm{d}r\bigg{\rVert}^{2}\right]+3\theta^{2}\mathbb{E}\left[\bigg{\lVert}\int_{t-\tau}^{t}(p_{r}+K^{T}Y_{r})\,\mathrm{d}r\bigg{\rVert}^{2}\right]$
$\qquad+6\mathbb{E}\left[\bigg{\lVert}\int_{t}^{s}\mathrm{d}W_{r}-\theta\int_{t-\tau}^{t}\mathrm{d}W_{r}\bigg{\rVert}^{2}\right]$
$\leq 3D_{4}((s-t)^{2}+\theta^{2}\tau^{2})+6d((s-t)+\theta^{2}\tau).$
Inserting the last inequality and (4.10) into (4.11) gives the error bound
$\mathbb{E}\left[\bigg{\lVert}\int_{t}^{t+\tau}(q_{s}-q_{t+\tau}+K(X_{s}-X_{t,\theta}))\,\mathrm{d}s\bigg{\rVert}^{2}\right]\leq 6L^{2}d(1+2\theta^{2})\tau^{3}+\left(2D_{4}L^{2}(1+3\theta^{2})+2D_{4}\omega_{f}^{-2}\right)\tau^{4}.$
Remark 4.7.
The assumption of Lemma 4.6 that $\lVert A_{\text{{PD}}}(Z_{t})\rVert^{2}$ is bounded in expectation is a standard assumption on the drift in Langevin sampling algorithms. In the case of overdamped Langevin diffusion, it is satisfied if the drift coefficient, i.e. $\nabla h$ in (3.2), is assumed to be Lipschitz-continuous, see Lemma 2 of [Dalalyan2017a].
In the case of primal-dual diffusion, this condition puts growth requirements on $\nabla g$ and $\nabla f^{\ast}$ instead. If, e.g., there are constants $\alpha,\beta$ such that $\lVert\nabla g(x)\rVert\leq\alpha+\beta\lVert x\rVert$ and $\lVert\nabla f^{\ast}(y)\rVert\leq\alpha+\beta\lVert y\rVert$ for all $x\in\mathcal{X}$ and $y\in\mathcal{Y}$, then the assumption is trivially satisfied. A further characterization of the stationary state $\pi_{\lambda}$ could help to replace this condition with more relaxed assumptions on $f$ and $g$.
The auxiliary inequality Lemma 4.5 and the bound on the discretization error in Lemma 4.6 allow us to give the following convergence result.
Theorem 4.8.
Let the assumptions of Lemma 4.6 be satisfied and assume further that $g,f^{\ast}$ are $\omega_{g}$- and $\omega_{f^{\ast}}$-strongly convex, respectively. Let the step sizes $\sigma,\tau$ and the extrapolation parameter $\theta$ be chosen such that $\theta\sigma\tau L^{2}\leq 1$ and
$\max\{(1+\omega_{g}\tau)^{-1},(1+\omega_{f^{\ast}}\sigma)^{-1}\}\leq\theta<1.$
Let the iterates $(X^{n},Y^{n})$ be generated by (1.4). Denoting $\mu^{n}$ the marginal distribution of $X^{n}$, we have
$W_{2}(\mu^{N},\mu_{\lambda})\leq\theta^{N/2}\mathcal{T}_{1,\bar{\lambda}^{-1}}(\pi^{0},\pi_{\lambda})+C_{2}(\tau),$
and further, if $\theta\sigma\tau L^{2}<1$ for the distribution $\pi^{n}$ of the joint variable $(X^{n},Y^{n})$ it holds
$\mathcal{T}_{1,\bar{\lambda}}(\pi^{N},\pi_{\lambda})\leq\frac{\theta^{N/2}}{(1-\theta\sigma\tau L^{2})^{1/2}}\mathcal{T}_{1,\bar{\lambda}^{-1}}(\pi^{0},\pi_{\lambda})+\frac{C_{2}(\tau)}{(1-\theta\sigma\tau L^{2})^{1/2}},$
with constants given by
$\bar{\lambda}\coloneqq\frac{1+\omega_{g}\tau}{1+\omega_{f^{\ast}}\sigma}\lambda,\quad C_{2}(\tau)\coloneqq\frac{(4dL_{\nabla g}^{2}\omega_{g}^{-1}+6dL^{2}(1+2\theta^{2})\omega_{f^{\ast}}^{-1})^{1/2}}{(1-\theta)^{1/2}}\tau+\mathcal{O}(\tau^{3/2}).$
Proof.
Let $(X^{n},Y^{n})$ be coupled with a process $\tilde{Z}_{t}=(\tilde{X}_{t},\tilde{Y}_{t})$, $t\geq 0$ by means of (4.6). Let $n\in\mathbb{N}_{0}$ be fixed and $t=n\tau$. Then
$\frac{2}{\tau}$ $\langle X^{n+1}-\tilde{X}_{t+\tau},\int_{t}^{t+\tau}(\tilde{p}_{s}-p^{\prime}+K^{T}(\tilde{Y}_{s}-\tilde{Y}_{t+\tau}))\,\mathrm{d}s\rangle$
$\leq\omega_{g}\lVert X^{n+1}-\tilde{X}_{t+\tau}\rVert^{2}+\frac{1}{\omega_{g}\tau^{2}}\bigg{\lVert}\int_{t}^{t+\tau}(\tilde{p}_{s}-p^{\prime}+K^{T}(\tilde{Y}_{s}-\tilde{Y}_{t+\tau}))\,\mathrm{d}s\bigg{\rVert}^{2},$
as well as
$\frac{2}{\tau}$ $\langle Y^{n+1}-\tilde{Y}_{t+\tau},\int_{t}^{t+\tau}(\tilde{q}_{s}-\tilde{q}_{t+\tau}-K(\tilde{X}_{s}-\tilde{X}_{t,\theta}))\,\mathrm{d}s\rangle$
$\leq\omega_{f^{\ast}}\lVert Y^{n+1}-\tilde{Y}_{t+\tau}\rVert^{2}+\frac{1}{\omega_{f^{\ast}}\tau^{2}}\bigg{\lVert}\int_{t}^{t+\tau}(\tilde{q}_{s}-\tilde{q}_{t+\tau}-K(\tilde{X}_{s}-\tilde{X}_{t,\theta}))\,\mathrm{d}s\bigg{\rVert}^{2}.$
We now employ the auxiliary inequality Lemma 4.5. Inserting the quadratic lower bounds for the symmetric Bregman distances and the two bounds above, we obtain the inequality
$\lVert X^{n+1}-\tilde{X}_{t+\tau}\rVert_{\tau^{-1}+\omega_{g}}^{2}+\lVert X^{n+1}-\tilde{X}_{t+\tau}-X^{n}+\tilde{X}_{t}\rVert_{\tau^{-1}}^{2}+\lVert Y^{n+1}-\tilde{Y}_{t+\tau}\rVert_{\sigma^{-1}+\omega_{f^{\ast}}}^{2}$
$\quad+2\langle K(X^{n+1}-\tilde{X}_{t+\tau}-X^{n}+\tilde{X}_{t}),Y^{n+1}-\tilde{Y}_{t+\tau}\rangle$
$\leq\lVert X^{n}-\tilde{X}_{t}\rVert_{\tau^{-1}}^{2}+\theta\lVert X^{n}-\tilde{X}_{t}-X^{n-1}+\tilde{X}_{t-\tau}\rVert_{\tau^{-1}}^{2}+\lVert Y^{n}-\tilde{Y}_{t}\rVert_{\sigma^{-1}}^{2}$
$\quad+2\theta\langle K(X^{n}-\tilde{X}_{t}-X^{n-1}+\tilde{X}_{t-\tau}),Y^{n}-\tilde{Y}_{t}\rangle$
$\quad+\underbrace{\frac{1}{\omega_{g}\tau^{2}}\bigg{\lVert}\int_{t}^{t+\tau}(\tilde{p}_{s}-p^{\prime}+K^{T}(\tilde{Y}_{s}-\tilde{Y}_{t+\tau}))\,\mathrm{d}s\bigg{\rVert}^{2}}_{\eqqcolon R_{n}^{(1)}}+\underbrace{\frac{1}{\omega_{f^{\ast}}\tau^{2}}\bigg{\lVert}\int_{t}^{t+\tau}(\tilde{q}_{s}-\tilde{q}_{t+\tau}-K(\tilde{X}_{s}-\tilde{X}_{t,\theta}))\,\mathrm{d}s\bigg{\rVert}^{2}}_{\eqqcolon R_{n}^{(2)}},$
We now define
$\Delta^{n}$ $\coloneqq\lVert X^{n}-\tilde{X}_{t}\rVert_{\tau^{-1}+\omega_{g}}^{2}+\lVert X^{n}-\tilde{X}_{t}-X^{n-1}+\tilde{X}_{t-\tau}\rVert_{\tau^{-1}}^{2}+\lVert Y^{n}-\tilde{Y}_{t}\rVert_{\sigma^{-1}+\omega_{f^{\ast}}}^{2}$
$\quad+2\langle K(X^{n}-\tilde{X}_{t}-X^{n-1}+\tilde{X}_{t-\tau}),Y^{n}-\tilde{Y}_{t}\rangle.$
Given the assumptions on $\theta$ and the above inequality, it holds
$\Delta^{n+1}$ $\leq\lVert X^{n}-\tilde{X}_{t}\rVert_{\tau^{-1}}^{2}+\theta\lVert X^{n}-\tilde{X}_{t}-X^{n-1}+\tilde{X}_{t-\tau}\rVert_{\tau^{-1}}^{2}+\lVert Y^{n}-\tilde{Y}_{t}\rVert_{\sigma^{-1}}^{2}$
$\quad+2\theta\langle K(X^{n}-\tilde{X}_{t}-X^{n-1}+\tilde{X}_{t-\tau}),Y^{n}-\tilde{Y}_{t}\rangle+R_{n}^{(1)}+R_{n}^{(2)}$
$\leq\theta\Delta^{n}+R_{n}^{(1)}+R_{n}^{(2)}$
Iterating this inequality over $n=0,\dots,N-1$ we obtain
$\Delta^{N}\leq\theta^{N}\Delta^{0}+\sum_{n=0}^{N-1}\theta^{n}(R_{N-n-1}^{(1)}+R_{N-n-1}^{(2)})$
Denote $T=N\tau$. By lower bounding the remaining inner product in $\Delta^{N}$, this implies
$\lVert X^{N}-\tilde{X}_{T}\rVert_{\tau^{-1}+\omega_{g}}^{2}+(1-\theta\tau\sigma L^{2})\lVert Y^{N}-\tilde{Y}_{T}\rVert_{\sigma^{-1}+\omega_{f^{\ast}}}^{2}$
$\quad\leq\theta^{N}\left(\lVert X^{0}-\tilde{X}^{0}\rVert_{\tau^{-1}+\omega_{g}}^{2}+\lVert Y^{0}-\tilde{Y}^{0}\rVert_{\sigma^{-1}+\omega_{f^{\ast}}}^{2}\right)+\sum_{n=0}^{N-1}\theta^{n}(R_{N-n-1}^{(1)}+R_{N-n-1}^{(2)})$
We now take expectations on both sides. The left side can be lower bounded by either $(\tau^{-1}+\omega_{g})\mathcal{W}_{2}^{2}(\mu^{N},\mu_{\lambda})$ or, in the joint variable, by $(\tau^{-1}+\omega_{g})(1-\theta\tau\sigma L^{2})\mathcal{T}^{2}_{1,\bar{\lambda}^{-1}}(\pi^{N},\pi_{\lambda})$. Since the initialization of the coupling was so far arbitrary (we only used that the marginal measures are $\pi_{0}$ and $\pi_{\lambda}$), we can choose the one that realizes the transport distance $\mathcal{T}_{1,\bar{\lambda}^{-1}}$ and obtain
$\mathbb{E}\left[\lVert X^{0}-\tilde{X}^{0}\rVert_{\tau^{-1}+\omega_{g}}^{2}+\lVert Y^{0}-\tilde{Y}^{0}\rVert_{\sigma^{-1}+\omega_{f^{\ast}}}^{2}\right]=(\tau^{-1}+\omega_{g})\mathcal{T}^{2}_{1,\bar{\lambda}^{-1}}(\pi^{0},\pi_{\lambda}).$
Furthermore, by Lemma 4.6 we know that for all $n=0,\dots,N-1$ it holds
$\mathbb{E}\left[R_{N-n-1}^{(1)}+R_{N-n-1}^{(2)}\right]\leq\underbrace{(4L_{\nabla g}^{2}\omega_{g}^{-1}+6L^{2}(1+2\theta^{2})\omega_{f^{\ast}}^{-1})}_{\eqqcolon M}d\tau+\mathcal{O}(\tau^{2})$
This means we can use $\sum_{n=0}^{N-1}\theta^{n}\leq(1-\theta)^{-1}$ and after dividing by $\tau^{-1}+\omega_{g}$ we obtain the two estimates
$\mathcal{W}_{2}^{2}(\mu^{N},\mu_{\lambda})\leq\theta^{N}\mathcal{T}^{2}_{1,\bar{\lambda}^{-1}}(\pi^{0},\pi_{\lambda})+\frac{Md}{1-\theta}\tau^{2}+\mathcal{O}(\tau^{3})$
and
$\mathcal{T}_{1,\bar{\lambda}^{-1}}^{2}(\pi^{N},\pi_{\lambda})\leq\frac{\theta^{N}}{1-\theta\sigma\tau L^{2}}\mathcal{T}^{2}_{1,\bar{\lambda}^{-1}}(\pi^{0},\pi_{\lambda})+\frac{Md}{(1-\theta\sigma\tau L^{2})(1-\theta)}\tau^{2}+\mathcal{O}(\tau^{3}),$
where we used that for $(\tau+\mathcal{O}(\tau^{2}))/(\tau^{-1}+\omega_{g})=\tau^{2}+\mathcal{O}(\tau^{3})$ for $\tau>0$. The statement of the theorem hence follows by taking the square root and using $\sqrt{a+b}\leq\sqrt{a}+\sqrt{b}$. ∎
In the last theorem, in the limit $N\to\infty$, the discrete time algorithm converges to its stationary distribution $\pi_{\lambda,\tau}$. From that, we immediately obtain the following bound between the invariant measures in discrete and continuous time.
Corollary 4.9.
Let the assumptions of Theorem 4.8 be satisfied. Then the following bounds hold:
$\mathcal{W}_{2}(\mu_{\lambda,\tau},\mu_{\lambda})$ $\leq C_{2}(\tau)+\mathcal{O}(\tau^{3/2}),$
$\mathcal{T}_{1,\bar{\lambda}^{-1}}(\pi_{\lambda,\tau},\pi_{\lambda})$ $\leq\frac{C_{2}(\tau)}{(1-\theta\sigma\tau L^{2})^{1/2}}+\mathcal{O}(\tau^{3/2}).$
The result Theorem 4.8 provides a bound between the samples produced by the algorithm and the corresponding stationary distribution in continuous time. While this result is typical in the Langevin algorithms literature, we here have to combine this with Theorem 3.11 in order to get a relation between samples and target distribution. The following result closes this final gap.
Corollary 4.10.
Let the assumptions of Theorems 3.11 and 4.8 be satisfied. Let $\pi^{n}$ be the distiribution of samples $(X^{n},Y^{n})$ generated by (1.4). Then the $x$-marginal $\mu^{n}$ of $\pi^{n}$ obeys
$\mathcal{W}_{2}(\mu^{N},\mu^{\ast})\leq\theta^{N/2}\mathcal{T}_{1,\bar{\lambda}^{-1}}(\pi^{0},\pi_{\lambda})+C_{1}(\lambda)+C_{2}(\tau),$
with the constants $C_{1},C_{2}$ from Theorems 3.11 and 4.8, respectively, which decay asymptotically like
$C_{1}(\lambda)$ $=D_{5}\lambda^{-1/2}+\mathcal{O}(\lambda^{-3/2}),\quad\textrm{as }\lambda\to\infty,$
$C_{2}(\tau)$ $=D_{6}\tau+\mathcal{O}(\tau^{3/2}),\quad\textrm{as }\tau\to 0.$
with constants gathered in $D_{5},D_{6}$. It immediately follows that $\mathcal{W}_{2}(\mu_{\lambda,\tau},\mu^{\ast})\leq C_{1}(\lambda)+C_{2}(\tau)$.
Corollary 4.11.
Let $(X^{n},Y^{n})$ be generated by the modified iteration (4.5), where the stochastic term $\xi$ is added in the primal variable before applying the proximal mapping. Under the same assumptions, the bounds of Theorem 4.8 and Corollaries 4.9 and 4.10 hold unchanged.
Proof.
As in the proof of Corollary 4.4, we have to modify the proof of the auxiliary inequality Lemma 4.5. It holds $p^{n+1}\in\partial g(X^{n+1})$, and accordingly we now choose $p^{\prime}\in\partial g(\tilde{X}_{t+\tau})$. Then, we start at the symmetric Bregman distance
$D_{g,\text{sym}}^{p^{n+1},p^{\prime}}(X^{n+1},\tilde{X}_{t+\tau})=\langle U^{n+1},p^{n+1}-p^{\prime}\rangle.$
The subsequent computations in the proof of Lemma 4.5 remain unchanged.
In order to bound the discretization error, the statement of Lemma 4.6 has to be changed to consider $p^{\prime}\in\partial g(X_{t+\tau})$ (instead of $p^{\prime}\in\partial g(X_{t+\tau}-\sqrt{2\tau}\xi)$). The error bound remains unchanged, the only difference in the proof appears in (4.9), where the term under the integral is now not $\mathbb{E}\left[\lVert X_{s}-X_{t+\tau}+\sqrt{2\tau}\xi\rVert^{2}\right]$, but $\mathbb{E}\left[\lVert X_{s}-X_{t+\tau}\rVert^{2}\right]$. Carrying out the subsequent computations in the same way, we arrive again at the upper bound $4L_{\nabla g}^{2}d\tau^{3}+\mathcal{O}(\tau^{4})$ for the primal discretization error. ∎
5 Numerical Examples
In the following, we demonstrate our results on the primal-dual Langevin sampling scheme on several numerical examples, first small-scale ones to demonstrate the convergence behaviour in detail and subsequently larger imaging inverse problems. We always simulate (2.6) in a time interval $[0,T]$ by running the discretized algorithm (1.4) with small step sizes $\tau,\sigma$ for $T/\tau$ steps. The step sizes are coupled via $\lambda\tau=\sigma$ and are chosen small enough so that $\tau\sigma L^{2}\leq 1$, i.e. the assumptions of Theorem 4.1 are satisfied and the continuous-time dynamics are approximated well.
5.1 Revisiting the Quadratic Case
We start by illustrating the toy example of two one-dimensional quadratic potentials that was considered in Theorems 3.5 and 3.6. Recall that the target distribution has density $\exp(-h(x))/Z$ with the potential $h$ defined as
$h(x)=f(kx)+g(x)=\frac{k^{2}}{2c_{f}}x^{2}+\frac{1}{2c_{g}}x^{2}=\frac{c_{g}k^{2}+c_{f}}{2c_{f}c_{g}}x^{2},$
hence the target distribution is Gaussian $\mathrm{N}(0,c_{f}c_{g}(c_{g}k^{2}+c_{f})^{-1})$. For the numerical simulation, we choose $c_{f}=1,c_{g}=2,k=1.5$. We simulate the primal-dual Langevin dynamics (2.5) for different values of $\lambda$ by running (1.4) (1.4). We set $\theta=1$ and set the step sizes $\tau$ and $\sigma$ via the relations $\sigma/\tau=\lambda$ and $\sigma\tau k^{2}=c$ where $c\leq 1$ is required for stability. We set $c=10^{-4}$ in order to approximate closely the continuous time dynamics, resulting in step sizes $\tau=10^{-2}\sqrt{\lambda}k$ and $\sigma=10^{-2}\sqrt{\lambda^{-1}}k$. Figure 3 shows the joint distribution of primal and dual samples $(X^{n},Y^{n})$ produced by (1.4). As $\lambda$ grows, the joint distribution of primal and dual variable concentrates along the linear subspace $y=kx/c_{f}$ and the primal marginal distribution converges to the target.
Figure 3: Approximation of the target $\mu^{\ast}=\exp(-h)$ for a one-dimensional quadratic $h$ by (1.4) and its continuous time diffusion limit (2.5). From left to right, results for $\lambda=1$, $\lambda=10$ and $\lambda=100$ are displayed. The center scatter plot shows pairs $(X^{n},Y^{n})$, the background color indicates the mass of the concentrated joint target $\pi^{\ast}$ (zero is blue, the mass is concentrated along the white line $y=kx/c_{f}$). The marginal boxes show the densities of the primal and dual targets $\mu^{\ast}$ and $\nu^{\ast}=(\partial f\circ K)_{\#}\mu^{\ast}$ (dashed red), the marginals $\mu^{\lambda},\nu^{\lambda}$ of the invariant solution of primal-dual diffusion (black solid) and samples drawn using (1.4) (grey histograms). For increasing $\lambda$, the joint distribution of samples tends to concentrate and the primal marginal converges to the target, as proven in Theorem 3.11.
5.2 TV Denoising from Gaussian Noise
For the potential $f\circ K$ being is dualized in the algorithm, we now consider the total variation (TV) functional, a typical model-based prior in imaging inverse problems. The (isotropic) TV functional is defined by
$\operatorname{TV}(x):=\sum_{i,j}\sqrt{(\nabla_{h}x)_{i,j}^{2}+(\nabla_{v}x)_{i,j}^{2}},$
where $\nabla_{h},\nabla_{v}$ denote horizontal and vertical finite differences in the pixelated image $x$. We consider a Gibbs prior on $x$ with density $\exp(-\alpha\operatorname{TV}(x))$ where $\alpha$ is a fixed regularization parameter (Note that this does only defines a probability distribution on a subspace orthogonal to constant functions since TV has a one-dimensional null space. With a log-likelihood that grows fast enough at infinity along this null space, the posterior is however well-defined.). The corresponding potential term can be written in the form $\alpha\operatorname{TV}(x)=\alpha\lVert Kx\rVert_{2,1}=f(Kx)$ where $K=(\nabla_{h},\nabla_{v})$ and $f=\alpha\lVert\cdot\rVert_{2,1}$ denotes a constant multiple of the nested $l_{2}$-$l_{1}$-norm (inner $l_{2}$ over vertical/horizontal component, outer $l_{1}$ over all pixels). We further assume that we make a noisy observation $\tilde{x}=x+\varepsilon$ of $x$, where the noise is distributed like $\varepsilon\sim\mathrm{N}(0,\sigma_{\varepsilon}^{2}I)$ for some observation variance $\sigma_{\varepsilon}^{2}$. The corresponding negative log-likelihood is
$g(x):=l_{\tilde{x}}(x)=\frac{1}{2\sigma_{\varepsilon}^{2}}\lVert x-\tilde{x}\rVert_{2}^{2},$
which is $\sigma^{-2}$-strongly convex. We now want to sample from the posterior $\mu^{\ast}$, which by Bayes law has density proportional to $\exp(-\alpha\lVert Kx\rVert_{2,1}-l_{\tilde{x}}(x))=\exp(-f(Kx)-g(x))$.
5.2.1 Toy Example in Two Dimensions
We start with a small-scale example in which we are able to compute accurate approximations of Wasserstein distances between the samples’ distribution and the true posterior. Consider the artificial case where the ‘image’ consists of $2\times 1$ pixels, i.e. $\mathcal{X}=\mathbb{R}^{2}$. Then the operator $K$ reduces to $Kx=x_{2}-x_{1}\in\mathbb{R}=\mathcal{Y}$ with $\lVert K\rVert=\sqrt{2}$, and $f$ is given by $f(y)=\alpha\left|y\right|$. The convex conjugate $f^{\ast}$ is the indicator function of the interval $[-\alpha,\alpha]$ and the proximal mapping $\operatorname{prox}_{\sigma f^{\ast}}$ is, for any $\sigma>0$, the projection onto $[-\alpha,\alpha]$.
The distribution $\mu_{n}$ of iterate $X^{n}$ in (1.4) is approximated by running finitely many, here $10^{4}$, independent Markov chains in parallel and using the resulting empirical distribution $\hat{\mu}_{n}$. In Figure 4, we show scatter plots of the samples at time $T$, where $T$ is chosen large enough that the empirical measures $\hat{\mu}_{n}$ remain stationary both qualitatively and in Wasserstein distance to the target (see below). As comparison, we show samples drawn using the algorithm Prox-Sub proposed in [Habring2023]. Prox-Sub solves the optimality condition $y\in\partial f(Kx)$ exactly at every step (by assuming that the subgradient can be evaluated), and hence corresponds exactly to the concentrated limit $\lambda=\infty$ of (1.4) explored in Section 3.3. The authors of [Habring2023] showed the convergence of the distribution of Prox-Sub samples to the target in 2-Wasserstein distance, total variation distance and Kullback-Leibler divergence, where the bias decreases depends as usual on the step size of the algorithm.
Figure 4: Scatter plots of $\hat{\mu}_{n}$ in the two-dimensional TV-denoising experiment. From left to right, the plots show approximations of the same posterior with values $\lambda=10^{1},10^{2},10^{3}$ and ‘$\lambda=\infty$’ (Prox-Sub). Horizontal and vertical axes correspond to the two coordinates in the primal variable $x\in\mathbb{R}^{2}$, the dual variable $y$ is not displayed. Grey dots show the samples at time $T$ produced by $10^{4}$ parallel chains, the black arrows indicate the samples’ mean and the principal components of their covariance (i.e. eigenvectors of the covariance matrix scaled in length by the corresponding eigenvalues). For reference, the blue arrows are the principal components of the posterior that should be approximated. The noisy observation $\tilde{x}$ is marked by a yellow triangle and the background color is a contour plot of the posterior density $\mu^{\ast}(x)$.
Using numerical integration, in this low-dimensional example we are able to compute an estimate of the normalization constant $Z=\int_{\mathcal{X}}\exp(-f(Kx)-g(x))\,\mathrm{d}x$. Using this, we can access an approximation of the posterior density $\mu^{\ast}(x)$. This allows us to compute approximations of the Wasserstein distances of current iterates to the target $\mathcal{W}_{2}(\hat{\mu}_{n},\mu^{\ast})$ without the need to rely on a second sampler. We evaluate the distance at 20 equidistantly spaced points in $[0,T]$ (i.e. for $\hat{\mu}_{kN},k=1,\dots,20$ where $N=T/20\tau$). The distances are computed using the Python optimal transport module [Flamary2021] and are displayed in Figure 5. As expected, the samples drawn by the Prox-Sub algorithm are closest to the target distribution (their bias in Wasserstein distance is bounded by a constant multiple of the step size $\tau$). For finite $\lambda$, the distributions $\hat{\mu}_{n}$ exhibit an additional strong bias to the target due to the dualization, which becomes smaller as $\lambda$ grows.
Figure 5: Wasserstein distances between the empirical measures $\hat{\mu}_{n}$ and the target $\mu^{\ast}$ in the two-dimensional TV-denoising experiment. We compare three different values of $\lambda$ and the algorithm Prox-Sub [Habring2023] corresponding formally to $\lambda=\infty$. The discretization via partial dualization of $f\circ K$ introduces a bias between samples and target that is significantly larger than the pure discretization bias of Prox-Sub. As $\lambda$ grows, the iteration approximates the purely primal setting and the bias is reduced.
5.2.2 Image Denoising
We now sample from the posterior of an image denoising task. Due to the higher dimension, we can not compute accurate approximations of the Wasserstein distances to the target anymore. We therefore resort to illustrating the behaviour of the algorithm (1.4) using moment estimates. A ground truth image normalized to intensities $[0,1]$ is distorted by noise with standard deviation $0.25$ (PSNR 12.0212.0212.02\,12.02dB in the noisy observation). We then draw $10^{6}$ samples with a single chain and compute the sample mean and pixel-wise variance after a burn-in phase of $31e$ 3 samples. The experiment is repeated three times for values $\lambda=1,10,100$, each with the same step size $\tau$. A fourth chain is computed using Prox-Sub with the same $\tau$. MMSE and pixel-wise variance estimates are reported in Figure 6. The sample mean (MMSE) estimates produced by the different chains converge quickly and are very similar for all chains (PSNR values between 23.8223.8223.82\,23.82dB and 23.8523.8523.85\,23.85dB). The pixel-wise variance, however, shows the same effect of overdispersion for smaller levels of $\lambda$ as in the lower-dimensional example.
While the ‘primal’ samples $X^{n}$ are overdispersed for finite $\lambda$, the dual iterates $Y^{n}$ characteristically exhibit underdispersion (see Figure 7). This effect was consistent throughout all our numerical experiments (compare also Figures 3 and 4). A heuristic explanation for overdispersion in the primal and underdispersion in the dual variable lies in the fact that the step in the dual variable solves the optimality condition $y\in\partial f(Kx)$ only incompletely. On the particle level in continuous time, when the primal variable $X_{t}$ is driven towards the tail regions of its marginal distribution by the Brownian motion, the dual variable is not evolved entirely to the corresponding position $\partial f(KX_{t})$ in the tails of its marginal, but rather remains near the mode of its marginal. The effect decreased with growing parameters $\lambda$ in all our experiments.
Figure 6: MMSE estimates and pixel-wise variance computed via the posterior sample mean in the $l_{2}$-TV denoising experiment. Top row shows the MMSE estimates, bottom row the corresponding pixel-wise (logarithmically scaled) variances. From left to right results for $\lambda=1,10,100$ and ‘$\lambda=\infty$’ (using Prox-Sub) are shown.
Figure 7: Pixel-wise (logarithmic) standard deviation of the dual variables $Y^{n}$, where an $l_{2}$-norm over horizontal and vertical component was applied. From left to right, we show the results for $\lambda=1,10,100$ and ‘$\lambda=\infty$’ (Prox-Sub). For the latter case, the algorithm computes the dual variable as $Y^{n}\in\partial f(KX^{n})$. For finite $\lambda$, the dual samples suffer from underdispersion, the effect vanishes as $\lambda$ grows and the dual values concentrate around the dual optimality condition.
5.3 Total Generalized Variation Based Image Denoising
In this experiment, we consider an example with an image prior based on second order total generalized variation (TGV). Like the TV functional, total generalized variation allows for discontinuous edges in images, but it generalizes TV in that it also employs higher-order derivatives. Acting as a penalty also on higher order image gradients, TGV-based MAP estimates avoid the typical staircasing effects of TV in smooth but non-constant regions of images. For optimization purposes, the second order TGV functional is usually defined as
$\operatorname{TGV}^{2}_{\alpha}(u)=\inf_{v\in V}J^{2}_{\alpha}(u,v),$ (5.1)
$J^{2}_{\alpha}(u,v)=\alpha_{1}\lVert\nabla u-v\rVert_{2,1}+\alpha_{0}\lVert Ev\rVert_{2,1}.$
Here $u\in\mathcal{U}$ is an image while $v$ is in the corresponding space $\mathcal{V}=\mathcal{U}\times\mathcal{U}$ of (discretized) gradients of $u$. $\alpha_{0},\alpha_{1}>0$ are regularization parameters. $\nabla=(\nabla_{h},\nabla_{v}):\mathcal{U}\to\mathcal{V}$ is a (forward discretized) gradient operator as in the case of TV, and $E:\mathcal{V}\to\mathcal{W}$ is a symmetrized (backward discretized) gradient operator that acts as second order derivative. We refer to Section 2.1 of [Bredies2015a] for details on the implementation of the gradient operators. In TGV-based variational minimization, the state space is typically expanded from the image space $\mathcal{U}$ to the joint space $\mathcal{X}=\mathcal{U}\times\mathcal{V}$ and the minimization over $u$ (in the variational problem) and $v$ (in 5.1) is carried out simultaneously. By applying primal-dual optimization methods, the iteration can be carried out efficiently without the need to evaluate the infimum in (5.1) explicitly [Bredies2015a]. In the same way, we want to avoid evaluating $\operatorname{TGV}^{2}_{\alpha}$ or its gradients explicitly in a sampling procedure. In order to recover the image $u$, we therefore propose to sample from a posterior of the form
$\mu^{\ast}(x)\propto p(\tilde{x}|x)\exp\left(-J^{2}_{\alpha}(u,v)\right),\qquad x=(u,v)\in\mathcal{U}\times\mathcal{V}=\mathcal{X},$
where $p(\tilde{x}|x)$ is the likelihood, for denosing with normally distributed noise given as before by
$-\log p(\tilde{x}|x)=\frac{1}{2\sigma_{\varepsilon}^{2}}\lVert u-\tilde{u}\rVert_{\mathcal{U}}^{2},\qquad x=(u,v),\tilde{x}=(\tilde{u},\tilde{v})\in\mathcal{X}$
We note that this is a relaxed version of regularization with $\operatorname{TGV}_{\alpha}^{2}$, since we dropped the requirement that $v$ solves (5.1). Practically, the prior $\exp(-J^{2}_{\alpha}(u,v))$ seems to yield very similar smoothing behaviour to the effect of total generalized variation, we leave its detailed analysis to future work. Note that the chosen target log-density is in the necessary form (1.1): $g(x)$ is, as before, the negative log-likelihood $\lVert u-\tilde{u}\rVert^{2}_{\mathcal{U}}/(2\sigma_{\varepsilon}^{2})$ for all $x=(u,v)$. We define the linear operator
$K=\begin{pmatrix}\nabla&-I\\ 0&E\end{pmatrix}:\mathcal{X}\to\mathcal{V}\times\mathcal{W}\eqqcolon\mathcal{Y},$
where $I$ is the identity on $\mathcal{V}$. The functional $f:\mathcal{Y}\to\mathbb{R}$ is given by $f(y)=\alpha_{1}\lVert v\rVert_{2,1}+\alpha_{0}\lVert w\rVert_{2,1}$ for all $y=(v,w)$. Due to the separable structure in $v$ and $w$, the proximal mapping of the convex conjugate $f^{\ast}$ is simply a componentwise projection in the dual $l_{2}$-$l_{\infty}$-norms in the respective spaces $\mathcal{V}$, $\mathcal{W}$.
We run the experiment with empirically set values of $\lambda$ that make sure the variance is not strongly overdispersed while keeping the computational expense moderate – recall that as $\lambda$ grows, $\tau$ needs to be chosen smaller, which slows down the convergence. $10^{5}$ samples are drawn in each run, we report the resulting sample mean and pixel-wise variance in Figure 8.
Figure 8: Results from the TGV-based denoising experiment. Left: noisy observation $\tilde{u}$, Center: MMSE estimate via sample mean, Right: Log-scaled pixel-wise standard deviation. Top is a moderately noisy phantom designed for TGV-regularization ($\sigma_{\varepsilon}=0.1$, PSNR of noisy image 19.9819.9819.98\,19.98dB, MMSE PSNR 34.0034.0034.00\,34.00dB), bottom a more strongly distorted, photorealistic image ($\sigma_{\varepsilon}=0.25$, PSNR of noisy image 12.0212.0212.02\,12.02dB, MMSE PSNR 20.0820.0820.08\,20.08dB).
6 Conclusion and Open Problems
We have analyzed a recently proposed algorithm for sampling from log-concave, non-smooth potentials which makes use of alternating primal-dual steps as is well-known in convex non-smooth optimization. We showed that the sampling algorithm corresponds to a stochastic differential equation and formulated the mean field limit Fokker-Planck equation. Using a simple coupling argument, contraction of the continuous-time limit to a stationary solution was proved. Unlike standard overdamped Langevin diffusion, the primal-dual Fokker-Planck equation is not a gradient flow in the corresponding Wasserstein space and does not have the target distribution as its stationary solution. A proposed correction of the Fokker-Planck equation forcing the target to be its stationary solution is possible, but eventually not useful for numerical implementation. However, the target distribution can be related to the stationary solution of the uncorrected scheme in the limit of the ratio of primal and dual step sizes growing to infinity. For the unmodified scheme, we made use of coupling techniques to prove a bound on Wasserstein distance between the stationary distribution and the target for a finite step size ratio.
With similar coupling arguments in discrete time, estimates in the typical spirit of convergence analysis of the optimization algorithm can be transferred to the sampling case. This allowed us to show stability of the sampling method in Wasserstein distance, resulting in the existence of a stationary measure in discrete time under strong convexity assumptions. Under the additional assumption of Lipschitz-smoothness of one of the potentials, we can bound the discretization error using a coupling between discrete and continuous time. This final bound implies a convergence result that relates the generated samples with the target. An important open question is whether the regularity assumptions on the potentials $f,g$ in the stability result and the discretization error bounds can be relaxed to not require differentiability.
Numerically, the results suggest that if there is a way to compute the subgradient of the dualized potential term, the additional bias due to the finite step size ratio makes the primal-dual samples consistently over-dispersed compared to discretizations of overdamped Langevin diffusion like the method of [Habring2023]. We currently see the primal-dual method’s use cases in two settings. If the subgradient of the dualized term is not available, but the proximal mapping of its convex conjugate is, the method is up to our knowledge the first applicable algorithm able to draw samples from an approximation of the posterior. Secondly, if the additional bias due to a larger finite step size ratio does not matter (because a correction step is added or because downstream tasks can be carried out with biased samples), the method can be applied with larger step sizes than a purely primal method, allowing for faster mixing time. To find out whether this makes the primal-dual method more efficient in applications, further numerical studies. Particularly, the combination of a proposal step using the primal-dual algorithm and a Metropolis-Hastings correction step seems like a promising method and should be analyzed further in future work.
Acknowledgements
We thank Jan Modersitzki (Lübeck) for fruitful discussions and predicting non-concentration of the dual variable.
[…]